### v4

# Patterson–Sullivan measures for relatively Anosov groups

Abstract.
We establish existence, uniqueness and ergodicity results for Patterson–Sullivan measures for relatively Anosov groups. As applications we obtain an entropy gap theorem and a strict concavity result for entropies associated to linear functionals.
Canary was partially supported by grant DMS-1906441 from the National Science Foundation. Zhang was partially supported by the NUS-MOE grant A-8000458-00-00. Zimmer was partially supported by a Sloan research fellowship and grants DMS-2105580 and DMS-2104381 from the National Science Foundation.
1. Introduction
Anosov subgroups and relatively Anosov subgroups of semisimple Lie groups are respectively natural generalizations of convex cocompact subgroups and geometrically finite subgroups of rank one semisimple Lie groups to the higher rank setting. Patterson–Sullivan measures for Anosov subgroups have been extensively studied, see [18, 38, 12, 33]. They have also been studied for relatively Anosov subgroups of the projective general linear group which preserve a properly convex domain, see [4, 5, 9, 42, 10]. More generally, Patterson–Sullivan measures can be constructed for transverse subgroups, of which Anosov subgroups, relatively Anosov subgroups, and discrete subgroups of rank one semisimple Lie groups are examples, see [15, 30, 31]. For a more detailed historical discussion of Patterson–Sullivan measures for discrete subgroups of semisimple Lie groups see [15, Sec. 1.1].
In this paper we study Patterson–Sullivan measures for relatively Anosov subgroups of semisimple Lie groups. The key new result here is that the Poincaré series associated to such a Patterson–Sullivan measure always diverges at its critical exponent if this critical exponent is finite. As a consequence, we establish uniqueness and ergodicity results for such Patterson–Sullivan measures. We then derive an entropy gap theorem and a strict concavity result for the entropy.
Delaying precise definitions until Sections 23 and 4, we informally introduce the notation necessary to state our main results. For the entire paper, $\mathsf{G}$ will be a connected semisimple Lie group without compact factors and with finite center. We fix a Cartan decomposition $\mathfrak{g}=\mathfrak{k}+\mathfrak{p}$ of the Lie algebra of $\mathsf{G}$, a Cartan subspace ${\mathfrak{a}}\subset{\mathfrak{p}}$, and a Weyl chamber $\mathfrak{a}^{+}\subset\mathfrak{a}$. Let $\Delta\subset{\mathfrak{a}}^{*}$ be the corresponding system of simple restricted roots, and let $\kappa:\mathsf{G}\to{\mathfrak{a}}^{+}$ denote the associated Cartan projection. Given a subset $\theta\subset\Delta$, we let $\operatorname{\mathsf{P}}_{\theta}\subset\mathsf{G}$ denote the associated parabolic subgroup and let $\operatorname{\mathcal{F}}_{\theta}=\mathsf{G}/\operatorname{\mathsf{P}}_{\theta}$ denote the associated flag manifold. We will always assume that $\theta$ is symmetric.
Suppose $\Gamma\subset\mathsf{G}$ is a non-elementary discrete subgroup which, as an abstract group, is relatively hyperbolic with respect to a finite collection $\mathcal{P}$ of subgroups in $\Gamma$. Let $\partial(\Gamma,\mathcal{P})$ denote its associated Bowditch boundary. Then $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$ if $\alpha\circ\kappa$ is proper on $\Gamma$ for all $\alpha\in\theta$, its limit set $\Lambda_{\theta}(\Gamma)$ in $\operatorname{\mathcal{F}}_{\theta}$ consists of mutually transverse flags and there exists a continuous $\Gamma$-equivariant map
$\xi:\partial(\Gamma,\mathcal{P})\rightarrow\operatorname{\mathcal{F}}_{\theta}$
which is a homeomorphism onto $\Lambda_{\theta}(\Gamma)$.
The action of $\mathsf{G}$ on $\operatorname{\mathcal{F}}_{\theta}$ preserves a vector valued cocycle $B_{\theta}:\mathsf{G}\times\operatorname{\mathcal{F}}_{\theta}\rightarrow{\mathfrak{a}}_{\theta}$, called the partial Iwasawa cocycle, whose image lies in a subspace ${\mathfrak{a}}_{\theta}\subset{\mathfrak{a}}$ associated to $\theta$. If $\phi\in{\mathfrak{a}}_{\theta}^{*}$, then the composition $\phi\circ B_{\theta}$ is a real valued cocycle, which allows us to define Patterson–Sullivan measures associated to such $\phi$. More precisely, if $\phi\in{\mathfrak{a}}_{\theta}^{*}$ and $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$, then a $\phi$-Patterson–Sullivan measure for $\Gamma$ of dimension $\beta$ is a Borel probability measure $\mu$ on $\operatorname{\mathcal{F}}_{\theta}$ where
  1. (1)
    ${\rm supp}(\mu)\subset\Lambda_{\theta}(\Gamma)$,
  2. (2)
    for every $\gamma\in\Gamma$ the measures $\gamma_{*}\mu$, $\mu$ are absolutely continuous and
    $\frac{d\gamma_{*}\mu}{d\mu}(F)=e^{-\beta\phi(B_{\theta}(\gamma^{-1},F))}$
    for $\mu$-almost every $F\in\operatorname{\mathcal{F}}_{\theta}$.
In the case when $\mathsf{G}=\mathsf{SO}_{0}(d,1)$ is the group of orientation-preserving isometries of real hyperbolic $d$-space $\operatorname{\mathbb{H}}^{d}$, there is a single simple restricted root $\Delta=\{\alpha\}$ and $\operatorname{\mathcal{F}}_{\alpha}$ naturally identifies with the geodesic boundary of $\operatorname{\mathbb{H}}^{d}_{\operatorname{\mathbb{R}}}$. Further (up to scaling), $\alpha\circ B_{\alpha}$ identifies with the usual Busemann cocycle. Since a discrete subgroup $\Gamma$ of $\mathsf{SO}_{0}(d,1)$ is relatively $\operatorname{\mathsf{P}}_{\alpha}$-Anosov if and only if it is geometrically finite, the above definition encompasses the classical notion of Patterson–Sullivan measures for geometrically finite Kleinian groups.
As in the classical theory, there is an associated Poincaré series and critical exponent. Let $\kappa_{\theta}:\mathsf{G}\rightarrow{\mathfrak{a}}_{\theta}$ denote the partial Cartan projection defined in Section 2. Then given $\phi\in{\mathfrak{a}}_{\theta}^{*}$, $s>0$ and a discrete group $\Gamma\subset\mathsf{G}$, the $\phi$-Poincaré series is
$Q_{\Gamma}^{\phi}(s):=\sum_{\gamma\in\Gamma}e^{-s\phi(\kappa_{\theta}(\gamma))}.$
The $\phi$-critical exponent, denoted $\delta^{\phi}(\Gamma)\in[0,+\infty]$, is the critical exponent of the above series, that is $Q_{\Gamma}^{\phi}(s)$ converges when $s>\delta^{\phi}(\Gamma)$ and diverges when $0<s<\delta^{\phi}(\Gamma)$. Given $\phi\in{\mathfrak{a}}_{\theta}^{*}$ let $\bar{\phi}\in{\mathfrak{a}}_{\theta}^{*}$ be the unique functional where $\bar{\phi}(\kappa_{\theta}(g))=\phi(\kappa_{\theta}(g^{-1}))$ for all $g\in\mathsf{G}$. Notice that $Q^{\phi}_{\Gamma}=Q^{\bar{\phi}}_{\Gamma}$ and so $\delta^{\phi}(\Gamma)=\delta^{\bar{\phi}}(\Gamma)$.
The main result of this paper is that if $\Gamma$ is relatively Anosov and its $\phi$-critical exponent is finite, then its $\phi$-Poincaré series diverges at its critical exponent.
Theorem 1.1 (Theorem 8.1).
If $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in{\mathfrak{a}}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$, then $Q_{\Gamma}^{\phi}$ diverges at its critical exponent.
Remark 1.2.
As mentioned above, in this paper we assume that relatively Anosov groups are non-elementary. Theorem 8.1 holds for elementary transverse groups (see Remark 8.5), however the non-elementary assumption is necessary for many of the applications of the theorem.
Later, we will discuss some important consequences of Theorem 1.1, and also provide an outline of its proof (in Sections 1.1 and 1.2 respectively). Notice that Theorem 1.1 fails in the setting of transverse groups, since the Poincaré series of any finitely generated, geometrically infinite discrete subgroup of $\mathsf{SO}(3,1)$ whose limit set is not all of $\partial\mathbb{H}^{3}$ converges at its critical exponent (see [13, Cor. 4.2] and [40, Thm. 2.17,  Cor. 2.18]).
We also provide a characterization of the functionals with finite critical exponent, surprisingly the only requirement is that $\phi(\kappa_{\theta}(\cdot))$ converges to infinity along escaping sequences in the group. We also show that if $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$, then $\phi(\kappa_{\theta}(\cdot))$ has linear lower and upper bounds in terms of the distance ${\rm d}_{M}$ on the Riemannian symmetric space $M$ associated to $\mathsf{G}$.
Theorem 1.3 (see Section 10).
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$ and $\phi\in\mathfrak{a}_{\theta}^{*}$. The following are equivalent:
  1. (1)
    $\lim_{n\rightarrow\infty}\phi(\kappa_{\theta}(\gamma_{n}))=+\infty$ for every sequence of distinct elements $\{\gamma_{n}\}$ in $\Gamma$.
  2. (2)
    $\delta^{\phi}(\Gamma)<+\infty$.
Moreover, if $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$, then the above conditions are equivalent to:
  1. (3)
    If $x_{0}\in M$ , there exist constants $c\geq 1$ and $C>0$ such that
    $\frac{1}{c}{\rm d}_{M}(\gamma(x_{0}),x_{0})-C\leq\phi(\kappa_{\theta}(\gamma))\leq c{\rm d}_{M}(\gamma(x_{0}),x_{0})+C$
    for all $\gamma\in\Gamma$.
We will observe later, see Section 1.3, that in Theorem 1.3 , Condition (3) is not equivalent to Conditions (1) and (2) without the assumption that $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$. Also, in the spirit of Sambarino’s analogous result for Anosov groups [38], we observe that Theorem 1.3 implies that $\phi\in\mathfrak{a}_{\theta}^{*}$ has finite critical exponent if and only if $\phi$ is positive on the $\theta$-Benoist limit cone, see Section 10.
1.1. Consequences of Theorem 1.1
We now recall some results from our previous work [15] from which we can derive several consequences of Theorem 1.1.
A discrete subgroup $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse if $\alpha\circ\kappa$ is proper on $\Gamma$ for all $\alpha\in\theta$ and its limit set $\Lambda_{\theta}(\Gamma)$ in $\operatorname{\mathsf{P}}_{\theta}$ consists of mutually transverse flags. A $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup acts on this limit set as a convergence group, so one can define the set of conical limit points $\Lambda^{\rm con}_{\theta}(\Gamma)\subset\Lambda_{\theta}(\Gamma)$ using the standard convergence group action definition. In this setting, we established in [15] the following analogue of the Hopf-Tsuji-Sullivan dichotomy.
Theorem 1.4 ([15, Thm. 1.4]).
Suppose $\Gamma\subset\mathsf{G}$ is a non-elementary $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup, $\phi\in\mathfrak{a}^{*}_{\theta}$ and $\delta:=\delta^{\phi}(\Gamma)<+\infty$.
  • If $Q_{\Gamma}^{\phi}(\delta)=+\infty$, then there exists a unique $\phi$-Patterson–Sullivan measure $\mu_{\phi}$ for $\Gamma$ of dimension $\delta$ and there exists a unique $\bar{\phi}$-Patterson–Sullivan measure $\mu_{\bar{\phi}}$ for $\Gamma$ of dimension $\delta$. Moreover:
    1. (1)
      $\mu_{\phi}(\Lambda_{\theta}^{\rm con}(\Gamma))=1=\mu_{\bar{\phi}}(\Lambda_{\theta}^{\rm con}(\Gamma))$.
    2. (2)
      The action of $\Gamma$ on $(\Lambda_{\theta}(\Gamma),\mu_{\phi})$ and $(\Lambda_{\theta}(\Gamma),\mu_{\bar{\phi}})$ is ergodic.
    3. (3)
      The action of $\Gamma$ on $(\Lambda_{\theta}(\Gamma)^{2},\mu_{\bar{\phi}}\otimes\mu_{\phi})$ is ergodic.
  • If $Q_{\Gamma}^{\phi}(\delta)<+\infty$, then $\mu(\Lambda_{\theta}^{\rm con}(\Gamma))=0$ for any $\phi$-Patterson–Sullivan measure $\mu$ for $\Gamma$.
If $\Gamma$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, then $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse. Thus, we obtain, as a consequence of Theorem 1.1 and Theorem 1.4, ergodicity and uniqueness results for their Patterson–Sullivan measures. These generalize results earlier obtained for Anosov groups (see [18, 38, 12, 33]).
Corollary 1.5.
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in{\mathfrak{a}}_{\theta}^{*}$ and $\delta:=\delta^{\phi}(\Gamma)<+\infty$. Then:
  1. (1)
    There is a unique $\phi$-Patterson–Sullivan measure $\mu_{\phi}$ for $\Gamma$ of dimension $\delta$ and a unique $\bar{\phi}$-Patterson–Sullivan measure $\mu_{\bar{\phi}}$ for $\Gamma$ of dimension $\delta$.
  2. (2)
    $\Gamma$ acts ergodically on $(\Lambda_{\theta}(\Gamma),\mu_{\phi})$ and $(\Lambda_{\theta}(\Gamma),\mu_{\bar{\phi}})$.
  3. (3)
    $\Gamma$ acts ergodically on $(\Lambda_{\theta}(\Gamma)\times\Lambda_{\theta}(\Gamma),\mu_{\phi}\otimes\mu_{\bar{\phi}})$.
In [15], we also established a criterion for when the critical exponent of a subgroup of a transverse group is strictly less than that of the entire group.
Theorem 1.6 ([15, Thm. 4.1]).
Suppose $\Gamma\subset\mathsf{G}$ is a non-elementary $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup, $\phi\in\mathfrak{a}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$. If $\Gamma_{0}$ is a subgroup of $\Gamma$ such that $Q_{\Gamma_{0}}^{\phi}$ diverges at its critical exponent and $\Lambda_{\theta}(\Gamma_{0})$ is a proper subset of $\Lambda_{\theta}(\Gamma)$, then $\delta^{\phi}(\Gamma)>\delta^{\phi}(\Gamma_{0}).$
Suppose that $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$. A subgroup $\Gamma_{0}\subset\Gamma$ is relatively quasiconvex if its action on its limit set $\Lambda(\Gamma_{0})\subset\partial(\Gamma,\mathcal{P})$ is geometrically finite. We prove that non-elementary, relatively quasiconvex subgroups of $\Gamma$ are themselves $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$ (see Proposition 9.4), and that the limit set of any infinite index quasiconvex subgroup of $\Gamma$ is a closed, proper subset of the limit set of $\Gamma$ (see Lemma 9.3). Combining these facts with Theorems 1.1 and 1.6, we may then prove the following result for relatively quasiconvex subgroups of infinite index.
Corollary 1.7 (Corollary 9.1).
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, and $\phi\in{\mathfrak{a}}_{\theta}^{*}$ such that $\delta^{\phi}(\Gamma)<+\infty$. If $\Gamma_{0}$ is an infinite index relatively quasiconvex subgroup of $(\Gamma,\mathcal{P})$, then
$\delta^{\phi}(\Gamma)>\delta^{\phi}(\Gamma_{0}).$
Finally, in [15], we established that the critical exponent is strictly concave on the space of linear functionals which diverge at their critical exponent, except when there is agreement of length functionals. More precisely, given $\phi\in{\mathfrak{a}}_{\theta}^{*}$ and $g\in\mathsf{G}$, the $\phi$-length of $g$ is
$\ell^{\phi}(g):=\lim_{n\rightarrow\infty}\frac{1}{n}\phi(\kappa_{\theta}(g^{n})).$
Theorem 1.8 ([15, Thm. 1.5]).
Suppose $\Gamma$ is a non-elementary $\mathsf{P}_{\theta}$-transverse subgroup of $\mathsf{G}$ and $\phi_{1},\phi_{2}\in\mathfrak{a}_{\theta}^{*}$ satisfy $\delta^{\phi_{1}}(\Gamma)=\delta^{\phi_{2}}(\Gamma)=1$. If $\phi=\lambda\phi_{1}+(1-\lambda)\phi_{2}$ for some $\lambda\in(0,1)$, then
$\delta:=\delta^{\phi}(\Gamma)\leq 1.$
Moreover, if $Q_{\Gamma}^{\phi}(\delta)=+\infty$, then equality occurs if and only if $\ell^{\phi_{1}}(\gamma)=\ell^{\phi_{2}}(\gamma)$ for all $\gamma\in\Gamma$.
Together, Theorem 1.1 and Theorem 1.8 give the following result.
Corollary 1.9.
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$ and $\phi_{1},\phi_{2}\in{\mathfrak{a}}_{\theta}^{*}$ satisfy $\delta^{\phi_{1}}(\Gamma)=\delta^{\phi_{2}}(\Gamma)=1$. If $\phi=\lambda\phi_{1}+(1-\lambda)\phi_{2}$ for some $\lambda\in(0,1)$, then
$\delta^{\phi}(\Gamma)\leq 1$
with equality if and only if $\ell^{\phi_{1}}(\gamma)=\ell^{\phi_{2}}(\gamma)$ for all $\gamma\in\Gamma$.
In the Zariski dense case, see [15, Cor. 1.6], a theorem of Benoist [2] implies that distinct length functions cannot agree.
Corollary 1.10.
Suppose $\Gamma\subset\mathsf{G}$ is a Zariski dense $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, and $\phi_{1},\phi_{2}\in{\mathfrak{a}}_{\theta}^{*}$ are distinct and satisfy $\delta^{\phi_{1}}(\Gamma)=\delta^{\phi_{2}}(\Gamma)=1$. If $\phi=\lambda\phi_{1}+(1-\lambda)\phi_{2}$ for some $\lambda\in(0,1)$, then $\delta^{\phi}(\Gamma)<1$.
1.2. Outline of the proof of Theorem 1.1
The strategy of the proof of Theorem 1.1 is inspired by earlier work of Blayac–Zhu [5] in the context of relatively hyperbolic groups preserving properly convex domains and Patterson–Sullivan measures defined using the Busemann functions associated to the Hilbert distance. The key technical result needed to prove Theorem 1.1 is that if $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$ and $\delta^{\phi}(\Gamma)<+\infty$, then the $\phi$-Poincaré series of any peripheral subgroup diverges at its critical exponent.
Theorem 1.11 (Theorem 7.1).
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in{\mathfrak{a}}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$. If $P\in\mathcal{P}$, then $Q_{P}^{\phi}$ diverges at its critical exponent.
Together, Theorems 1.6 and 1.11 imply that $\delta^{\phi}(P)<\delta^{\phi}(\Gamma)$ for all $P\in\mathcal{P}$. We may then adapt arguments of Dal’bo–Otal–Piegné [17] to our setting to conclude Theorem 1.1.
The proof of Theorem 1.11 makes use of Hironaka’s famous result on the resolution of singularities. More precisely, in Section 7, we relate the Poincaré series $Q_{P}^{\phi}(s)$ associated to a peripheral subgroup $P$ to an integral of the form
[…]
where $\lambda$ is the Lebesgue measure, $R_{1},\dots,R_{m}:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}$ are positive rational functions that are defined everywhere and $\ell_{1},\dots,\ell_{m}$ are real numbers such that $R_{1}^{\ell_{1}}\cdots R_{m}^{\ell_{m}}$ is a proper function. In Section 6, we use Hironaka’s resolution of singularities to show that the integral in Equation (1) diverges at its critical exponent. This in turn implies that the Poincaré series associated to any peripheral subgroup diverges at its critical exponent.
In Blayac and Zhu’s setting, the integral in Equation (1) involves a single rational function (i.e.  $m=1$) and one can deduce that it diverges at its critical exponent from a result of Benoist–Oh [3, Prop. 7.2] (see the proof of  [5, Lem. 8.9]). The case when $m>1$ is more technical and our use of Hironaka’s resolution of singularities to understand the integral at infinity is motivated by Benoist and Oh’s arguments.
1.3. Conditions (1) and (2) are not equivalent to (3) in Theorem 1.3
We now give an example which shows that Condition (3) in Theorem 1.3 may fail to be equivalent to Conditions (1) and (2) when $\operatorname{\mathsf{P}}_{\theta}$ contains a simple factor of $\mathsf{G}$.
Example 1.12.
Let $\mathsf{G}=\mathsf{PSL}(2,\operatorname{\mathbb{R}})\times\mathsf{PSL}(2,\operatorname{\mathbb{R}})$. Note that
$\mathfrak{a}:=\left\{{\rm diag}(a_{1},\dots,a_{4}):a_{1}+a_{2}=0=a_{3}+a_{4}\right\}$
is a Cartan subspace of $\mathsf{G}$ and $\Delta:=\{\alpha_{1},\alpha_{1}^{\prime}\}\subset\mathfrak{a}^{*}$ is a system of simple restricted roots, where
$\alpha_{1}({\rm diag}(a_{1},\dots,a_{4}))=a_{1}-a_{2}\quad\text{and}\quad\alpha_{1}^{\prime}({\rm diag}(a_{1},\dots,a_{4}))=a_{3}-a_{4}.$
Set $\theta:=\{\alpha_{1}\}$, and note that
${\mathfrak{a}}_{\theta}=\{{\rm diag}(a,-a,0,0):a\in\operatorname{\mathbb{R}}\}$
and
$\kappa_{\theta}((g,h))=(\log\sigma_{1}(g),\log\sigma_{2}(g),0,0)$
where $\sigma_{1}(g)\geq\sigma_{2}(g)$ are the singular values of some (any) lift of $g$ to $\operatorname{\mathsf{SL}}(2,\operatorname{\mathbb{R}})$. Also,
$\operatorname{\mathsf{P}}_{\theta}=\operatorname{\mathsf{P}}_{1}\times\mathsf{PSL}(2,\operatorname{\mathbb{R}})$
where $\operatorname{\mathsf{P}}_{1}\subset\mathsf{PSL}(2,\operatorname{\mathbb{R}})$ stabilizes the line $\operatorname{\mathbb{R}}\cdot\begin{pmatrix}1\\ 0\end{pmatrix}$. In particular, $\operatorname{\mathsf{P}}_{\theta}$ contains a simple factor of $\mathsf{G}$.
[…]
$\gamma:=(u,h)\in\mathsf{G}\quad\text{ and }\quad\eta:=(h,\operatorname{\mathrm{id}})\in\mathsf{G},$
then $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}:=\{\left\langle\gamma\right\rangle\}$. Furthermore,
$\lim_{n\rightarrow\infty}\alpha_{1}(\kappa_{\theta}(g_{n}))=+\infty$
for every sequence of distinct elements $\{g_{n}\}$ in $\Gamma$, so (1) in Theorem 1.3 holds.
On the other hand, the Riemannian symmetric space $M$ associated to $\mathsf{G}$ is the product
$M=\operatorname{\mathbb{H}}^{2}_{\operatorname{\mathbb{R}}}\times\operatorname{\mathbb{H}}^{2}_{\operatorname{\mathbb{R}}}$
where $\operatorname{\mathbb{H}}^{2}_{\operatorname{\mathbb{R}}}$ is the real hyperbolic 2-space. Then for any $p_{0}=(x_{0},y_{0})\in M$ we have
$\liminf_{n\rightarrow\infty}\frac{1}{n}\log{\rm d}_{M}(\gamma^{n}(p_{0}),p_{0})=\liminf_{n\rightarrow\infty}\frac{1}{n}\log{\rm d}_{\operatorname{\mathbb{H}}_{\operatorname{\mathbb{R}}}^{2}}(h^{n}(y_{0}),y_{0})>0$
since $h$ is hyperbolic and $u$ is unipotent. However,
$\limsup_{n\rightarrow\infty}\frac{1}{n}\log\phi(\kappa_{\theta}(\gamma^{n}))=\limsup_{n\rightarrow\infty}\frac{1}{n}\log\frac{\sigma_{1}(u^{n})}{\sigma_{2}(u^{n})}=0$
since $u$ is unipotent. So (3) in Theorem 1.3 does not hold.
Acknowledgements
We thank the referee for their comments and their careful reading of the original manuscript.
2. Semisimple Lie groups
In this expository section we introduce some of the notation used throughout the paper. All the notation introduced is the same as in [15].
As in the introduction (and the rest of the paper), let $\mathsf{G}$ be a connected semisimple Lie group without compact factors and with finite center. Fix a Cartan decomposition
${\mathfrak{g}}={\mathfrak{k}}\oplus{\mathfrak{p}}$
of the Lie algebra ${\mathfrak{g}}$ of $G$, a Cartan subspace $\mathfrak{a}\subset\mathfrak{p}$, and a Weyl chamber by ${\mathfrak{a}}^{+}\subset{\mathfrak{a}}$. Let $\Sigma^{+}$ be the set of positive restricted roots associated to $\mathfrak{a}^{+}$, let $\Delta\subset\mathfrak{a}^{*}$ be the system of simple restricted roots associated to $\mathfrak{a}^{+}$, and let $\operatorname{\mathsf{K}}\subset\mathsf{G}$ denote the maximal compact Lie subgroup whose Lie algebra is ${\mathfrak{k}}$.
2.1. Cartan projection
The Cartan projection is the map $\kappa:\mathsf{G}\rightarrow{\mathfrak{a}}^{+}$ with the defining property that $\kappa(g)\in{\mathfrak{a}}^{+}$ is the unique element such that
$g=me^{\kappa(g)}\ell$
for some $m,\ell\in\operatorname{\mathsf{K}}$ (in general $m$ and $\ell$ are not uniquely determined by $g$). Such a decomposition $g=me^{\kappa(g)}\ell$ is called a $\mathsf{KAK}$-decomposition of $g$, see [22, Chap. IX, Thm. 1.1].
There is a unique involutive linear map $\iota:{\mathfrak{a}}\rightarrow{\mathfrak{a}}$, called the opposition involution, such that $\iota(-{\mathfrak{a}}^{+})={\mathfrak{a}}^{+}$ and
$\iota(\kappa(g))=\kappa(g^{-1})$
for all $g\in\mathsf{G}$. The dual of $\iota$ satisfies
$\iota^{*}(\Delta)=\Delta.$
As in the introduction we let $M$ denote the standard Riemannian symmetric space associated to $\mathsf{G}$, and let ${\rm d}_{M}$ denote the distance function on $M$ induced by the Riemannian metric. As a $\mathsf{G}$-space, $M=\mathsf{G}/\operatorname{\mathsf{K}}$, and ${\rm d}_{M}$ satisfies
${\rm d}_{M}(g\operatorname{\mathsf{K}},h\operatorname{\mathsf{K}})=\left\|\kappa(g^{-1}h)\right\|\quad\text{for all}\quad g,h\in\mathsf{G},$
[…]
Lemma 2.1 (see e.g. [29, Lem. 2.3]).
If $g,h\in\mathsf{G}$, then
$\left\|\kappa(gh)-\kappa(h)\right\|\leq\left\|\kappa(g)\right\|\quad\text{and}\quad\left\|\kappa(gh)-\kappa(g)\right\|\leq\left\|\kappa(h)\right\|.$
2.2. Parabolic subgroups and flag manifolds
Given a subset $\theta\subset\Delta$, the parabolic subgroup associated to $\theta$, denoted by $\operatorname{\mathsf{P}}_{\theta}\subset\mathsf{G}$, is the normalizer of
${\mathfrak{u}}_{\theta}:=\bigoplus_{\alpha\in\Sigma^{+}_{\theta}}{\mathfrak{g}}_{\alpha}$
where $\Sigma^{+}_{\theta}:=\Sigma^{+}-{\rm Span}(\Delta-\theta)$. Then the flag manifold associated to $\theta$ is
$\operatorname{\mathcal{F}}_{\theta}:=\mathsf{G}/\mathsf{P}_{\theta}.$
In this paper will only consider the case when $\theta\subset\Delta$ is symmetric, that is $\iota^{*}(\theta)=\theta$. In this case, there is a unique open $\mathsf{G}$-orbit in
$\operatorname{\mathcal{F}}_{\theta}\times\operatorname{\mathcal{F}}_{\theta}$
and we say that two flags $F_{1},F_{2}\in\operatorname{\mathcal{F}}_{\theta}$ are transverse if $(F_{1},F_{2})$ is in this orbit. For a flag $F\in\operatorname{\mathcal{F}}_{\theta}$, we let
$\mathcal{Z}_{F}\subset\operatorname{\mathcal{F}}_{\theta}$
denote the set of flags that are not transverse to $F$.
Following the notation in [21], we define a map
$U_{\theta}:\mathsf{G}\rightarrow\operatorname{\mathcal{F}}_{\theta}$
by fixing a $\mathsf{KAK}$-decomposition $g=m_{g}e^{\kappa(g)}\ell_{g}$ for each $g\in\mathsf{G}$ and then letting $U_{\theta}(g):=m_{g}\operatorname{\mathsf{P}}_{\theta}$. One can show that if $\alpha(\kappa(g))>0$ for all $\alpha\in\theta$, then $U_{\theta}(g)$ is independent of the choice of $\mathsf{KAK}$-decomposition, see [22, Chap. IX, Thm. 1.1], and hence $U_{\theta}$ is continuous on the set
$\left\{g\in\mathsf{G}:\alpha(\kappa(g))>0\text{ for all }\alpha\in\theta\right\}.$
The action of $\mathsf{G}$ on $\operatorname{\mathcal{F}}_{\theta}$ has the following north-south like dynamics.
Proposition 2.2 (see e.g. [15, Prop. 2.3]).
Suppose $\theta\subset\Delta$ is symmetric, $F^{\pm}\in\operatorname{\mathcal{F}}_{\theta}$ and $\{g_{n}\}$ is a sequence in $\mathsf{G}$. The following are equivalent:
  1. (1)
    $U_{\theta}(g_{n})\rightarrow F^{+}$, $U_{\theta}(g_{n}^{-1})\rightarrow F^{-}$ and $\lim_{n\rightarrow\infty}\alpha(\kappa(g_{n}))=\infty$ for every $\alpha\in\theta$,
  2. (2)
    $g_{n}(F)\to F^{+}$ for all $F\in\operatorname{\mathcal{F}}_{\theta}-\mathcal{Z}_{F^{-}}$, and this convergence is uniform on compact subsets of $\operatorname{\mathcal{F}}_{\theta}-\mathcal{Z}_{F^{-}}$.
  3. (3)
    $g_{n}^{-1}(F)=F^{-}$ for all $F\in\operatorname{\mathcal{F}}_{\theta}-\mathcal{Z}_{F^{+}}$, and this convergence is uniform on compact subsets of $\operatorname{\mathcal{F}}_{\theta}-\mathcal{Z}_{F^{+}}$.
  4. (4)
    there are open sets $\mathcal{U}^{\pm}\subset\operatorname{\mathcal{F}}_{\theta}$ such that $g_{n}(F)\to F^{+}$ for all $F\in\mathcal{U}^{+}$ and $g_{n}^{-1}(F)\to F^{-}$ for all $F\in\mathcal{U}^{-}$.
2.3. Fundamental weights and partial Cartan projections
For any $\alpha\in\Delta$, let $\omega_{\alpha}\in{\mathfrak{a}}^{*}$ denote the associated fundamental weight.
Given a subset $\theta\subset\Delta$, the partial Cartan subspace associated to $\theta$ is
${\mathfrak{a}}_{\theta}:=\{H\in{\mathfrak{a}}:\alpha(H)=0\text{ for all }\alpha\in\Delta-\theta\}.$
Then there is a unique projection
$p_{\theta}:{\mathfrak{a}}\rightarrow{\mathfrak{a}}_{\theta}$
such that $\omega_{\alpha}(X)=\omega_{\alpha}(p_{\theta}(X))$ for all $\alpha\in\theta$ and $X\in\mathfrak{a}$. The partial Cartan projection associated to $\theta$ is
$\kappa_{\theta}:=p_{\theta}\circ\kappa:\mathsf{G}\to\mathfrak{a}_{\theta}.$
One can show that $\{\omega_{\alpha}|_{{\mathfrak{a}}_{\theta}}:\alpha\in\theta\}$ is a basis of ${\mathfrak{a}}_{\theta}^{*}$ and hence we will identify
${\mathfrak{a}}_{\theta}^{*}={\rm Span}\{\omega_{\alpha}:\alpha\in\theta\}\subset\mathfrak{a}^{*}.$
Note that $\omega_{\alpha}(\kappa_{\theta}(g))=\omega_{\alpha}(\kappa(g))$ for all $\alpha\in\theta$ and $g\in\mathsf{G}$ and so
$\phi(\kappa_{\theta}(g))=\phi(\kappa(g))$ (2)
for all $\phi\in{\mathfrak{a}}_{\theta}^{*}$ and $g\in\mathsf{G}$.
2.4. The partial Iwasawa cocycle
Let $\operatorname{\mathsf{U}}:=\exp({\mathfrak{u}}_{\Delta})$. The Iwasawa decomposition states that the map
$(k,a,u)\in\operatorname{\mathsf{K}}\times\exp({\mathfrak{a}})\times\operatorname{\mathsf{U}}\mapsto kau\in\mathsf{G}$
is a diffeomorphism, see [32, Chap. VI, Prop. 6.46]. Using this, Quint [37] defined the Iwasawa cocycle
$B:\mathsf{G}\times\operatorname{\mathcal{F}}_{\Delta}\rightarrow{\mathfrak{a}}$
with the defining property that $gk\in\operatorname{\mathsf{K}}\cdot\exp(B(g,F))\cdot\operatorname{\mathsf{U}}$ for all $(g,F)\in\mathsf{G}\times\operatorname{\mathcal{F}}_{\Delta}$, where $k\in\operatorname{\mathsf{K}}$ is an element such that $F=k\operatorname{\mathsf{P}}_{\Delta}$.
For any $\theta\subset\Delta$, note that $\operatorname{\mathsf{P}}_{\Delta}\subset\operatorname{\mathsf{P}}_{\theta}$, so the identity map on $\mathsf{G}$ induces a surjection $\Pi_{\theta}:\operatorname{\mathcal{F}}_{\Delta}\to\operatorname{\mathcal{F}}_{\theta}$. The partial Iwasawa cocycle is the map
$B_{\theta}:\mathsf{G}\times\operatorname{\mathcal{F}}_{\theta}\rightarrow{\mathfrak{a}}_{\theta}$
defined by $B_{\theta}(g,F)=p_{\theta}(B(g,F^{\prime}))$ for some (all) $F^{\prime}\in\Pi_{\theta}^{-1}(F)$. By [37, Lem. 6.1 and 6.2], this is a well-defined cocycle, that is
$B_{\theta}(gh,F)=B_{\theta}(g,hF)+B_{\theta}(h,F)$
for all $g,h\in\mathsf{G}$ and $F\in\operatorname{\mathcal{F}}_{\theta}$.
2.5. The Linear Case
We now briefly describe the above notations when $\mathsf{G}=\mathsf{PSL}(d,\operatorname{\mathbb{R}})$. Let $\mathfrak{sl}(d,\operatorname{\mathbb{R}})=\mathfrak{k}+\mathfrak{p}$ denote the standard Cartan decomposition of $\mathfrak{sl}(d,\operatorname{\mathbb{R}})$, that is
$\mathfrak{k}=\{X\in\mathfrak{sl}(d,\operatorname{\mathbb{R}}):{{}^{t}X}=-X\}\quad\text{and}\quad\mathfrak{p}=\{X\in\mathfrak{sl}_{d}(\operatorname{\mathbb{R}}):{{}^{t}X}=X\}.$
Also let
$\mathfrak{a}:=\{{\rm diag}(a_{1},\ldots,a_{d})\in\mathfrak{sl}(d,\operatorname{\mathbb{R}}):\ a_{1}+\cdots+a_{d}=0\}\subset\mathfrak{p}$
denote the standard Cartan subspace with the standard positive Weyl chamber
$\mathfrak{a}^{+}:=\{{\rm diag}(a_{1},\ldots,a_{d})\in\mathfrak{a}:\ a_{1}\geq\dots\geq a_{d}\}\subset\mathfrak{p}.$
Then the opposite involution is given by
$\iota({\rm diag}(a_{1},\ldots,a_{d}))={\rm diag}(-a_{d},\ldots,-a_{1})$
and the Cartan projection is given by
$\kappa(g)=(\log\sigma_{1}(g),\cdots,\log\sigma_{d}(g))$
where $\sigma_{1}(g)\geq\cdots\geq\sigma_{d}(g)$ are the singular values of any lift of $g$ to $\mathsf{SL}(d,\operatorname{\mathbb{R}})$.
The standard system of simple restricted roots is $\Delta:=\{\alpha_{1},\dots,\alpha_{d-1}\}\subset\mathfrak{a}^{*}$ where
$\alpha_{j}({\rm diag}(a_{1},\ldots,a_{d}))=a_{j}-a_{j+1}$
for all ${\rm diag}(a_{1},\ldots,a_{d})\in\mathfrak{a}$. Then the fundamental weights are given by
$\omega_{\alpha_{j}}({\rm diag}(a_{1},\ldots,a_{d}))=a_{1}+\dots+a_{j}$
and so $\iota^{*}(\omega_{\alpha_{j}})=\omega_{\alpha_{d-j}}$.
Let $e_{1},\dots,e_{d}$ denote the standard basis of $\operatorname{\mathbb{R}}^{d}$ and let $F_{0}^{j}:=\left\langle e_{1},\dots,e_{j}\right\rangle$ be the subspace spanned by $e_{1},\dots,e_{j}$. Then given $\theta=\{\alpha_{j_{1}},\dots,\alpha_{j_{n}}\}\subset\Delta$ with $j_{1}<j_{2}<\dots<j_{n}$, the parabolic subgroup $\operatorname{\mathsf{P}}_{\theta}\subset\mathsf{PSL}(d,\operatorname{\mathbb{R}})$ is the stabilizer of the flag
$(F_{0}^{j_{1}},F_{0}^{j_{2}},\cdots,F^{j_{n}}_{0})$
and $\operatorname{\mathcal{F}}_{\theta}$ is the partial flag manifold
$\mathcal{F}_{\theta}=\left\{(F^{j_{1}},\dots,F^{j_{n}}):\ \mathrm{dim}\left(F^{j_{i}}\right)=j_{i}\text{ and }F^{j_{1}}\subset F^{j_{2}}\subset\cdots\subset F^{j_{n}}\right\}.$
In the case when $\theta$ is symmetric, two flags $F_{1},F_{2}\in\mathcal{F}_{\theta}$ are transverse if and only if they are transverse in the usual sense: $F_{1}^{j}\oplus F_{2}^{d-j}=\operatorname{\mathbb{R}}^{d}$ for all $\alpha_{j}\in\theta$.
In the linear case we often replace subscripts of the form $\theta=\{\alpha_{j_{1}},\dots,\alpha_{j_{n}}\}$ with $j_{1},\cdots,j_{n}$. So for instance
$\operatorname{\mathcal{F}}_{1,d-1}=\operatorname{\mathcal{F}}_{\{\alpha_{1},\alpha_{d-1}\}}$
denotes the line/hyperplane partial flag manifold.
2.6. Properties of unipotent subgroups
Recall that a $d\times d$ real matrix $A$ is unipotent if 1 is the only eigenvalue of $A$ (over $\operatorname{\mathbb{C}}$) and a subgroup $\mathsf{U}\subset\mathsf{GL}(d,\operatorname{\mathbb{R}})$ is called unipotent if every element in $\mathsf{U}$ is unipotent. When $\mathsf{G}$ has trivial center, a subgroup $\mathsf{U}\subset\mathsf{G}$ is unipotent if ${\rm Ad}(\mathsf{U})\subset\operatorname{\mathsf{SL}}(\mathfrak{g})$ is unipotent. Given a Lie subgroup $\mathsf{H}\subset\mathsf{G}$, the unipotent radical of $\mathsf{H}$ is the maximal connected normal unipotent subgroup of $\mathsf{H}$.
Unipotent subgroups have the following well-known properties.
Proposition 2.3.
Suppose $Z(\mathsf{G})$ is trivial and $\mathfrak{u}$ is the Lie algebra of a connected unipotent subgroup $\mathsf{U}\subset\mathsf{G}$. Then:
  1. (1)
    The exponential map induces a diffeomorphism $\mathfrak{u}\rightarrow\mathsf{U}$.
  2. (2)
    Let $\lambda$ be a measure on $\mathfrak{u}$ which is obtained by pulling back the Lebesgue measure on $\operatorname{\mathbb{R}}^{\dim\mathfrak{u}}$ via a linear isomorphism (notice that any two measures obtained this way are scalar multiplies of each other). Then the push-forward $\mu:=\exp_{*}\lambda$ is a Haar measure on $\mathsf{U}$.
  3. (3)
    If $\Phi:\mathsf{G}\rightarrow\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ is a representation, then $\Phi(\mathsf{U})\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ is unipotent.
  4. (4)
    There exists $C>0$ such that
    $\left\|\kappa(\exp(Y))\right\|\leq C+C\log(1+\left\|Y\right\|)$
    for all $Y\in\mathfrak{u}$.
  5. (5)
    For any $\alpha\in\Delta$ there exist $M_{\alpha}\in\operatorname{\mathbb{N}}$, $C_{\alpha}>1$ and a positive everywhere defined rational function $R_{\alpha}:\mathfrak{u}\rightarrow\operatorname{\mathbb{R}}$ where
    $\frac{1}{C_{\alpha}}R_{\alpha}(Y)^{1/M_{\alpha}}\leq e^{\omega_{\alpha}(\kappa(\exp(Y)))}\leq C_{\alpha}R_{\alpha}(Y)^{1/M_{\alpha}}$
    for all $Y\in\mathfrak{u}$.
We provide a proof of Proposition 2.3 in Appendix A.
3. Relatively hyperbolic groups
In this expository section we recall one of the many equivalent definitions of a relatively hyperbolic group, for more background and other definitions see [8, 19, 35, 41].
3.1. Relatively hyperbolic groups
Recall that the action, by homeomorphisms, of a group $\Gamma$ on a compact metrizable space $M$ is said to be a (discrete) convergence group action if whenever $\{\gamma_{n}\}$ is a sequence of distinct elements in $\Gamma$, then there are points $x,y\in M$ and a subsequence $\{\gamma_{n_{j}}\}$ so that $\gamma_{n_{j}}(z)$ converges to $x$ for all $z\in M-\{y\}$ and the convergence is uniform on compact subsets of $M-\{y\}$.
Suppose that $\Gamma$ acts on $M$ as a convergence group, then:
  • A point $x\in M$ is a conical limit point if there exist $a,b\in M$ distinct and a sequence $\{\gamma_{n}\}$ in $\Gamma$ such that $\gamma_{n}(x)\to a$ and $\gamma_{n}(y)\to b$ for all $y\in M-\{x\}$.
  • An element $\gamma\in\Gamma$ is parabolic if it has infinite order and fixes exactly one point in $M$.
  • A infinite order subgroup $H\subset\Gamma$ is parabolic if it fixes some point of $M$ and each infinite order element in $H$ is parabolic. The fixed point of a parabolic subgroup is called a parabolic point.
  • A parabolic point $x\in M$ is bounded if the quotient ${\rm Stab}_{\Gamma}(x)\backslash(M-\{x\})$ is compact.
Finally, $\Gamma$ is called a geometrically finite convergence group if every point in $M$ is either a conical limit point or a bounded parabolic point.
Definition 3.1.
Given a finitely generated group $\Gamma$ and a collection $\mathcal{P}$ of finitely generated infinite subgroups, we say that $\Gamma$ is hyperbolic relative to $\mathcal{P}$, or that $(\Gamma,\mathcal{P})$ is relatively hyperbolic, if $\Gamma$ acts on a compact perfect metrizable space $M$ as a geometrically finite convergence group and the maximal parabolic subgroups are exactly the set
$\{\gamma P\gamma^{-1}:P\in\mathcal{P},\gamma\in\Gamma\}.$
To avoid confusion with parabolic subgroups in the Lie group sense, we also sometimes call a maximal parabolic subgroup of $\Gamma$ a peripheral subgroup.
Remark 3.2.
Notice that by definition we assume that a relatively hyperbolic group is non-elementary (i.e.  $M$ is perfect and hence is infinite) and finitely generated.
By a theorem of Bowditch [8, Thm. 9.4], given a relatively hyperbolic group $(\Gamma,\mathcal{P})$, any two compact perfect metrizable spaces satisfying Definition 3.1 are $\Gamma$-equivariantly homeomorphic. This unique topological space is then denoted by $\partial(\Gamma,\mathcal{P})$ and called the Bowditch boundary of $(\Gamma,\mathcal{P})$.
3.2. The Groves–Manning cusp space
Given a relatively hyperbolic group $(\Gamma,\mathcal{P})$, Groves and Manning [20] constructed a Gromov hyperbolic space which $\Gamma$ acts on proper discontinuously so that subgroups in $\mathcal{P}$ preserve “combinatorial horoballs.” These spaces, now called Groves–Manning cusp spaces, are central tools in the study of relatively hyperbolic groups. We recall their construction here.
Definition 3.3.
Suppose $Y$ is a graph with the simplicial distance ${\rm d}_{Y}$. The combinatorial horoball $\mathcal{H}(Y)$ is the graph, also equipped with the simplicial distance, that has vertex set $Y^{(0)}\times\operatorname{\mathbb{N}}$ and two types of edges:
  • vertical edges joining vertices $(v,n)$ and $(v,n+1)$,
  • horizontal edges joining vertices $(v,n)$ and $(w,n)$ when ${\rm d}_{Y}(v,w)\leq 2^{n-1}$.
Definition 3.4.
Suppose that $(\Gamma,\mathcal{P})$ is relatively hyperbolic. A finite symmetric generating set $S\subset\Gamma$ is adapted if $S\cap P$ is a generating set of $P$ for every $P\in\mathcal{P}$. Given such an $S$, we let ${\mathcal{C}}(\Gamma,S)$ and ${\mathcal{C}}(P,S\cap P)$ denote the associated Cayley graphs. Then the associated Groves–Manning cusp space, denoted ${\mathcal{C}}_{GM}(\Gamma,\mathcal{P},S)$, is obtained from the Cayley graph ${\mathcal{C}}(\Gamma,S)$ by attaching, for each $P\in\mathcal{P}$ and each coset $\gamma P\in\Gamma/P$, a copy of the combinatorial horoball $\mathcal{H}(\gamma{\mathcal{C}}(P,S\cap P))$ by identifying $\gamma{\mathcal{C}}(P,S\cap P)\subset{\mathcal{C}}(\Gamma,S)$ with the $n=1$ level of $\mathcal{H}(\gamma{\mathcal{C}}(P,S\cap P))$.
Theorem 3.5 (Groves–Manning [20, Thm. 3.25]).
If $(\Gamma,\mathcal{P})$ is relatively hyperbolic and $S$ is an adapted finite generating set, then ${\mathcal{C}}_{GM}(\Gamma,\mathcal{P},S)$ is a proper geodesic Gromov hyperbolic space such that
  1. (1)
    $\Gamma$ acts properly discontinuously on ${\mathcal{C}}_{GM}(\Gamma,\mathcal{P},S)$ by isometries,
  2. (2)
    every point in $X$ is within a uniformly bounded distance of a bi-infinite geodesic, and
  3. (3)
    there exists a $\Gamma$-equivariant homeomorphism between $\partial_{\infty}{\mathcal{C}}_{GM}(\Gamma,\mathcal{P},S)$, the Gromov boundary of ${\mathcal{C}}_{GM}(\Gamma,\mathcal{P},S)$, and $\partial(\Gamma,\mathcal{P})$.
4. Discrete subgroups of semisimple Lie groups
In this expository section we introduce three classes of discrete subgroups in $\mathsf{G}$ and state some of their basic properties. In the rest of the paper, we assume that $\theta\subset\Delta$ is symmetric.
4.1. Divergent groups
A discrete subgroup $\Gamma\subset\mathsf{G}$ is called $\operatorname{\mathsf{P}}_{\theta}$-divergent if
$\lim_{n\rightarrow+\infty}\min_{\alpha\in\theta}\alpha(\kappa_{\theta}(\gamma_{n}))=+\infty$
whenever $\{\gamma_{n}\}$ is a sequence of distinct elements in $\Gamma$. The limit set $\Lambda_{\theta}(\Gamma)$ of such a subgroup is the set of accumulation points of $\{U_{\theta}(\gamma):\gamma\in\Gamma\}$. We note that in the literature, divergent groups are sometimes called regular groups (e.g. [27]).
The limit set of a divergent group can be used to compactify it.
Lemma 4.1 (see e.g. [15, Prop. 2.3]).
If $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-divergent, then the set $\Gamma\cup\Lambda_{\theta}(\Gamma)$ has a topology that makes it a compactification of $\Gamma$. More precisely:
  1. (1)
    $\Gamma\cup\Lambda_{\theta}(\Gamma)$ is a compact metrizable space.
  2. (2)
    If $\Gamma$ has the discrete topology, then $\Gamma\hookrightarrow\Gamma\cup\Lambda_{\theta}(\Gamma)$ is an embedding.
  3. (3)
    If $\Lambda_{\theta}(\Gamma)$ has the subspace topology from $\operatorname{\mathcal{F}}_{\theta}$, then $\Lambda_{\theta}(\Gamma)\hookrightarrow\Gamma\cup\Lambda_{\theta}(\Gamma)$ is an embedding.
  4. (4)
    A sequence $\{\gamma_{n}\}$ in $\Gamma$ converges to $F$ in $\Lambda_{\theta}(\Gamma)$ if and only if
    $\lim_{n\rightarrow+\infty}\min_{\alpha\in\theta}\alpha(\kappa_{\theta}(\gamma_{n}))=+\infty\quad\text{and}\quad U_{\theta}(\gamma_{n})\rightarrow F.$
  5. (5)
    The natural left action of $\Gamma$ on $\Gamma\cup\Lambda_{\theta}(\Gamma)$ is by homeomorphisms.
4.2. Transverse groups
A $\operatorname{\mathsf{P}}_{\theta}$-divergent subgroup $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse if $\Lambda_{\theta}(\Gamma)$ is a transverse subset of $\operatorname{\mathcal{F}}_{\theta}$, i.e. distinct pairs of flags in $\Lambda_{\theta}(\Gamma)$ are transverse. We note that in the literature, transverse groups are sometimes called regular antipodal groups (e.g. [27]).
One crucial feature of $\operatorname{\mathsf{P}}_{\theta}$-transverse groups is that they act as a convergence group on their limit sets.
Proposition 4.2.
[28, Prop. 5.38] If $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse, then $\Gamma$ acts on $\Lambda_{\theta}(\Gamma)$ as a convergence group. In particular, if $\Gamma$ is non-elementary, then $\Gamma$ acts on $\Lambda_{\theta}(\Gamma)$ minimally, and $\Lambda_{\theta}(\Gamma)$ is perfect.
When $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse, the set of conical limit points for the action of $\Gamma$ on $\Lambda_{\theta}(\Gamma)$ is called the $\theta$-conical limit set and is denoted $\Lambda^{\rm con}_{\theta}(\Gamma)$.
4.3. Relatively Anosov subgroups
There are several equivalent definitions of relatively Anosov groups. The definition we use comes from [26].
A $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$, a finite collection of subgroups of $\Gamma$, if $(\Gamma,\mathcal{P})$ is relatively hyperbolic with Bowditch boundary $\partial(\Gamma,\mathcal{P})$ and there is a continuous $\Gamma$-equivariant map
$\xi\colon\partial(\Gamma,\mathcal{P})\to\operatorname{\mathcal{F}}_{\theta}$
which is a homeomorphism onto $\Lambda_{\theta}(\Gamma)$. Observe that such a $\xi$ is unique, so we refer to it as the limit map of $\Gamma$.
The next result shows that this limit maps plays nicely with the Gromov boundary of a Groves–Manning cusp space.
Proposition 4.3.
Suppose $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$, with limit map $\xi:\partial(\Gamma,\mathcal{P})\rightarrow\operatorname{\mathcal{F}}_{\theta}$. Let $X$ be a Groves–Manning cusp space for $(\Gamma,\mathcal{P})$, and let $b_{0}\in X$. If $\{\gamma_{n}\}$ is a sequence in $\Gamma$ and $\gamma_{n}(b_{0})\rightarrow x\in\partial_{\infty}X=\partial(\Gamma,\mathcal{P})$, then $U_{\theta}(\gamma_{n})\rightarrow\xi(x)$.
Proof.
Since $\operatorname{\mathcal{F}}_{\theta}$ is compact, it suffices to show that every convergent subsequence of $\{U_{\theta}(\gamma_{n})\}$ converges to $\xi(x)$. Suppose $U_{\theta}(\gamma_{n_{j}})\rightarrow F^{+}$. Passing to a further subsequence we can suppose that $\gamma_{n_{j}}^{-1}(b_{0})\rightarrow y\in\partial(\Gamma,\mathcal{P})$ and $U_{\theta}(\gamma_{n_{j}}^{-1})\rightarrow F^{-}$. Then, by properties of Gromov hyperbolic spaces, $\gamma_{n_{j}}(z)\rightarrow x$ for all $z\in\partial(\Gamma,\mathcal{P})-\{y\}$. Also, by Proposition 2.2,
$\lim_{j\rightarrow\infty}\gamma_{n_{j}}(F)=F^{+}$
for all $F\in\operatorname{\mathcal{F}}_{\theta}$ transverse to $F^{-}$.
Notice that Proposition 2.2 implies that $F^{-}\in\Lambda_{\theta}(\Gamma)$ and so $F^{-}=\xi(y^{\prime})$ for some $y^{\prime}\in\partial(\Gamma,\mathcal{P})$. Fix $z\in\partial(\Gamma,\mathcal{P})-\{y,y^{\prime}\}$. Then
$\xi(x)=\lim_{j\rightarrow\infty}\xi(\gamma_{n_{j}}(z))=\lim_{n\rightarrow\infty}\gamma_{n_{j}}\xi(z)=F^{+}$
since $\xi(z)$ is transverse to $F^{-}=\xi(y^{\prime})$. ∎
The following theorem was established in  [43] when $\mathsf{G}=\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$. In Appendix B we will explain why it is also true in the following setting. In the following theorem and elsewhere in the paper, given a Lie group $\mathsf{H}$ let $\mathsf{H}^{0}$ denote the connected component of the identity in $\mathsf{H}$.
Theorem 4.4.
Assume $Z(\mathsf{G})$ is trivial and $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$. Suppose $\Gamma\subset\mathsf{G}$ is a non-elementary $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$.
  1. (1)
    If $X$ is a Groves–Manning cusp space for $(\Gamma,\mathcal{P})$ and $M:=\mathsf{G}/\operatorname{\mathsf{K}}$ is a Riemannian symmetric space associated to $\mathsf{G}$, then there exist $c>1$, $C>0$ such that
    $\frac{1}{c}{\rm d}_{M}(\gamma\operatorname{\mathsf{K}},\operatorname{\mathsf{K}})-C\leq{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})\leq c{\rm d}_{M}(\gamma\operatorname{\mathsf{K}},\operatorname{\mathsf{K}})+C$
    for all $\gamma\in\Gamma$.
  2. (2)
    If $P\in\mathcal{P}$, then $P$ is a cocompact lattice in a closed Lie subgroup $\mathsf{H}\subset\mathsf{G}$ with finitely many components. Moreover,
    1. (a)
      $\mathsf{H}=\mathsf{L}\ltimes\mathsf{U}$ where $\mathsf{L}$ is compact and $\mathsf{U}$ is the unipotent radical of $\mathsf{H}$.
    2. (b)
      $\mathsf{H}^{0}=\mathsf{L}^{0}\times\mathsf{U}$ and $\mathsf{L}^{0}$ is Abelian.
Example 1.12 provides an example where $\operatorname{\mathsf{P}}_{\theta}$ contains a simple factor of $\mathsf{G}$ and the conclusions of Theorem 4.4 fail.
4.4. Helpful reductions
We first explain why one can often reduce to the case where the center $Z(\mathsf{G})$ of $\mathsf{G}$ is trivial and $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$. We then explain how one can often reduce to the case where $\mathsf{G}=\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$.
Decompose the Lie algebra ${\mathfrak{g}}$ of $G$ into a product of simple Lie algebras, ${\mathfrak{g}}=\oplus_{j=1}^{m}{\mathfrak{g}}_{j}$. For each $1\leq j\leq m$, let $\mathsf{G}_{j}\subset\mathsf{G}$ denote the closed connected normal subgroup with Lie algebra ${\mathfrak{g}}_{j}$. Then
$\mathsf{G}=\mathsf{G}_{1}\cdots\mathsf{G}_{m}$
is an almost direct product and $\mathsf{G}_{1},\dots,\mathsf{G}_{m}$ are called the simple factors of $\mathsf{G}$.
Proposition 4.5 ([15, Prop. 2.9]).
Suppose $\theta\subset\Delta$ is symmetric and $\mathsf{H}:=Z(\mathsf{G})\prod\{\mathsf{G}_{j}:\mathsf{G}_{j}\subset\operatorname{\mathsf{P}}_{\theta}\}$. Let $p:\mathsf{G}\rightarrow\mathsf{G}^{\prime}:=\mathsf{G}/\mathsf{H}$ be the quotient map. Then:
  1. (1)
    $\mathsf{G}^{\prime}$ is a semisimple Lie group without compact factors and with trivial center.
  2. (2)
    There is a Cartan decomposition ${\mathfrak{g}}^{\prime}={\mathfrak{k}}^{\prime}+{\mathfrak{p}}^{\prime}$ of the Lie algebra of $\mathsf{G}^{\prime}$, a Cartan subspace ${\mathfrak{a}}^{\prime}\subset{\mathfrak{p}}^{\prime}$, a system of simple restricted roots $\Delta^{\prime}\subset({\mathfrak{a}}^{\prime})^{*}$ and a subset $\theta^{\prime}\subset\Delta^{\prime}$ such that
    $p(\operatorname{\mathsf{P}}_{\theta})=\operatorname{\mathsf{P}}^{\prime}_{\theta^{\prime}}$
    (where $\operatorname{\mathsf{P}}^{\prime}_{\theta^{\prime}}$ is the parabolic subgroup of $\mathsf{G}^{\prime}$ associated to $\theta^{\prime}$). Moreover, $\operatorname{\mathsf{P}}_{\theta^{\prime}}^{\prime}$ contains no simple factors of $\mathsf{G}^{\prime}$.
  3. (3)
    ${\rm d}p$ induces an isomorphism of the partial Cartan subspaces ${\mathfrak{a}}_{\theta}$ and ${\mathfrak{a}}_{\theta^{\prime}}^{\prime}$. Moreover, the partial Cartan projections satisfies
    ${\rm d}p(\kappa_{\theta}(g))=\kappa_{\theta^{\prime}}^{\prime}(p(g))\quad\text{for all $g\in\mathsf{G}$}.$
  4. (4)
    The map $\xi:\operatorname{\mathcal{F}}_{\theta}\rightarrow\operatorname{\mathcal{F}}_{\theta^{\prime}}=\mathsf{G}^{\prime}/\operatorname{\mathsf{P}}_{\theta^{\prime}}^{\prime}$ defined by $\xi(g\operatorname{\mathsf{P}}_{\theta})=p(g)\operatorname{\mathsf{P}}_{\theta^{\prime}}^{\prime}$ is a diffeomorphism which preserves transversality. Moreover, the partial Iwasawa cocycles satisfies
    ${\rm d}p(B_{\theta}(g,F))=B^{\prime}_{\theta^{\prime}}(p(g),\xi(F))\quad\text{for all $g\in\mathsf{G}$ and $F\in\operatorname{\mathcal{F}}_{\theta}$}.$
Using the discussion in [21, Section 3] it is possible to prove the following result which allows one to reduce many calculations to the linear case, see  [15, Prop. B.1] for details. The statement of the result uses the notation introduced in Section 2.5.
Proposition 4.6.
For any symmetric $\theta\subset\Delta$ and $\chi\in\sum_{\alpha\in\theta}\operatorname{\mathbb{N}}\cdot\omega_{\alpha}$ there exist $d\in\operatorname{\mathbb{N}}$, an irreducible linear representation $\Phi:\mathsf{G}\rightarrow\mathsf{SL}(d,\operatorname{\mathbb{R}})$ and a $\Phi$-equivariant smooth embedding
$\xi:\operatorname{\mathcal{F}}_{\theta}\rightarrow\operatorname{\mathcal{F}}_{1,d-1}(\operatorname{\mathbb{R}}^{d})$
such that:
  1. (1)
    $F_{1},F_{2}\in\operatorname{\mathcal{F}}_{\theta}$ are transverse if and only if $\xi(F_{1})$ and $\xi(F_{2})$ are transverse.
  2. (2)
    There exists $N\in\operatorname{\mathbb{N}}$ such that
    $\log\sigma_{1}(\Phi(g))=N\chi(\kappa(g))$
    for all $g\in\mathsf{G}$.
  3. (3)
    $\alpha_{1}(\kappa(\Phi(g)))=\min_{\alpha\in\theta}\alpha(\kappa(g))$ for all $g\in\mathsf{G}$.
  4. (4)
    If $\min_{\alpha\in\theta}\alpha(\kappa(g))>0$, then
    $\xi(U_{\theta}(g))=U_{1,d-1}(\Phi(g)).$
  5. (5)
    $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-divergent (respectively $\operatorname{\mathsf{P}}_{\theta}$-transverse) if and only if $\Phi(\Gamma)$ is $\operatorname{\mathsf{P}}_{1,d-1}$-divergent (respectively $\operatorname{\mathsf{P}}_{1,d-1}$-transverse). Moreover, in this case
    $\xi(\Lambda_{\theta}(\Gamma))=\Lambda_{1,d-1}(\Phi(\Gamma)).$
  6. (6)
    If $\Gamma\subset\mathsf{G}$ is discrete and $\mathcal{P}$ is a finite collection of subgroups in $\Gamma$, then $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$ if and only if $\Phi(\Gamma)$ is $\operatorname{\mathsf{P}}_{1,d-1}$-Anosov relative to $\mathcal{P}^{\prime}:=\{\Phi(P):P\in\mathcal{P}\}$.
Remark 4.7.
Part (6) is not explicitly stated in  [15, Prop. B.1], however it follows immediately from part (5) and the definitions.
5. Multiplicative estimates
It is a general principle from linear algebra that if $U_{\theta}(A^{-1})$ is uniformly transverse to $U_{\theta}(B)$, then $\kappa_{\theta}(AB)$ is coarsely equal to $\kappa_{\theta}(A)+\kappa_{\theta}(B)$, see, for example, [6, Lem. A.7]. We make use of two manifestations of this principle, the first in the context of transverse groups and the second in the context of relatively Anosov groups. Previous instances of this principle in our work include [14, Lem. 6.2] and [15, Prop. 6.3].
[…]
Proposition 5.1.
Suppose $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse and ${\rm d}_{\operatorname{\mathcal{F}}_{\theta}}$ is a distance on $\operatorname{\mathcal{F}}_{\theta}$ which is induced by a Riemannian metric. For any $\epsilon>0$ there exists $C=C(\epsilon)>0$ such that: if $\gamma,\eta\in\Gamma$ and ${\rm d}_{\operatorname{\mathcal{F}}_{\theta}}(U_{\theta}(\gamma^{-1}),U_{\theta}(\eta))>\epsilon$, then
$\left\|\kappa_{\theta}(\gamma\eta)-\kappa_{\theta}(\gamma)-\kappa_{\theta}(\eta)\right\|\leq C.$
Proof.
We use the following special case of [6, Lem. A.7].
Lemma 5.2.
Let $e_{1},\dots,e_{d}$ denote the standard basis of $\operatorname{\mathbb{R}}^{d}$. Suppose $g_{1},g_{2}\in\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ have singular value decomposition $g_{1}=m_{1}a_{1}\ell_{1}$ and $g_{2}=m_{2}a_{2}\ell_{2}$. Then
$\sigma_{1}(g_{1})\sigma_{1}(g_{2})\sin(\theta)\leq\sigma_{1}(g_{1}g_{2})\leq\sigma_{1}(g_{1})\sigma_{1}(g_{2})$
where $\theta:=\angle\left(m_{2}\left\langle e_{1}\right\rangle,\ell_{1}^{-1}\left\langle e_{2},\dots,e_{d}\right\rangle\right)$ is the Euclidean angle between the subspaces $m_{2}\left\langle e_{1}\right\rangle$ and $\ell_{1}^{-1}\left\langle e_{2},\dots,e_{d}\right\rangle$.
If the proposition fails, then for every $n\geq 1$ there exist $\gamma_{n},\eta_{n}\in\Gamma$ where
${\rm d}_{\operatorname{\mathcal{F}}_{\theta}}(U_{\theta}(\gamma_{n}^{-1}),U_{\theta}(\eta_{n}))>\epsilon$
and
$\left\|\kappa_{\theta}(\gamma_{n}\eta_{n})-\kappa_{\theta}(\gamma_{n})-\kappa_{\theta}(\eta_{n})\right\|\geq n.$
Lemma 2.1 implies that $\{\gamma_{n}\}$ and $\{\eta_{n}\}$ are escaping sequences in $\Gamma$. So by passing to a subsequence we can suppose that $U_{\theta}(\gamma_{n}^{-1})\rightarrow F_{1}\in\Lambda_{\theta}(\Gamma)$ and $U_{\theta}(\eta_{n})\rightarrow F_{2}\in\Lambda_{\theta}(\Gamma)$. Then ${\rm d}_{\operatorname{\mathcal{F}}_{\theta}}(F_{1},F_{2})\geq\epsilon$ and hence $F_{1}$ and $F_{2}$ are transverse (since $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse).
Since $\{\omega_{\alpha}|_{\mathfrak{a}_{\theta}}:\alpha\in\theta\}$ is a basis for $\mathfrak{a}_{\theta}^{*}$, after passing to a subsequence there exists some $\chi\in\sum_{\alpha\in\theta}\operatorname{\mathbb{N}}\cdot\omega_{\alpha}$ such that
$\lim_{n\rightarrow\infty}\left|\chi\Big{(}\kappa_{\theta}(\gamma_{n}\eta_{n})-\kappa_{\theta}(\gamma_{n})-\kappa_{\theta}(\eta_{n})\Big{)}\right|=\infty.$
Let $N\in\operatorname{\mathbb{N}}$, $\Phi:\mathsf{G}\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{R}})$ and $\xi:\operatorname{\mathcal{F}}_{\theta}\rightarrow\operatorname{\mathcal{F}}_{1,d-1}(\operatorname{\mathbb{R}}^{d})$ satisfy Proposition 4.6 for $\chi$. Then by Proposition 4.6 Claim (2),
$\log\sigma_{1}(\Phi(g))=N\chi(\kappa(g))=N\chi(\kappa_{\theta}(g))$
for all $g\in\mathsf{G}$.
Let $\hat{\gamma}_{n}:=\Phi(\gamma_{n})$ and $\hat{\eta}_{n}:=\Phi(\eta_{n})$. Since $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-transverse, $\alpha(\kappa(\gamma_{n}))\rightarrow+\infty$ for all $\alpha\in\theta$. So by Proposition 4.6 Claim (4),
$\lim_{n\rightarrow\infty}U_{1,d-1}(\hat{\gamma}_{n}^{-1})=\lim_{n\rightarrow\infty}\xi(U_{\theta}(\gamma_{n}^{-1}))=\xi(F_{1}).$
Likewise, $U_{1,d-1}(\hat{\eta}_{n})\rightarrow\xi(F_{2})$. Since $F_{1}$ and $F_{2}$ are transverse, Proposition 4.6 Claim (1) implies that $\xi(F_{1})$ and $\xi(F_{2})$ are transverse. So Lemma 5.2 implies that there exists a constant $C>0$ such that
$\left|\log\sigma_{1}(\hat{\gamma}_{n}\hat{\eta}_{n})-\log\sigma_{1}(\hat{\gamma}_{n})-\log\sigma_{1}(\hat{\eta}_{n})\right|\leq C$
for all $n\geq 1$. So,
$\left|\chi\Big{(}\kappa_{\theta}(\gamma_{n}\eta_{n})-\kappa_{\theta}(\gamma_{n})-\kappa_{\theta}(\eta_{n})\Big{)}\right|\leq\frac{C}{N}$
and we have a contradiction. ∎
Our result for relatively Anosov groups involves a choice of Groves–Manning cusp space.
Proposition 5.3.
Suppose $\Gamma\subset\mathsf{G}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov relative to $\mathcal{P}$ and $X$ is a Groves–Manning cusp space for $(\Gamma,\mathcal{P})$. There exists $C>0$ such that: if $f:[0,T]\rightarrow X$ is a geodesic with $f(0)=\operatorname{\mathrm{id}}$ and $f(T)\in\Gamma$, then
$\left\|\kappa_{\theta}(f(T))-\kappa_{\theta}(f(t))-\kappa_{\theta}(f(t)^{-1}f(T))\right\|\leq C$
whenever $t\in[0,T]$ and $f(t)\in\Gamma$.
Proof.
Suppose not. Then for every $n\geq 1$ there is a geodesic $f_{n}:[0,T_{n}]\rightarrow X$ and some $t_{n}\in[0,T_{n}]$ such that $f_{n}(0)=\operatorname{\mathrm{id}}$, $f_{n}(t_{n})\in\Gamma$, $f_{n}(T_{n})\in\Gamma$, and
$\left\|\kappa_{\theta}(f_{n}(T_{n}))-\kappa_{\theta}(f_{n}(t_{n}))-\kappa_{\theta}(f_{n}(t_{n})^{-1}f_{n}(T_{n}))\right\|\geq n.$
Let $\gamma_{n}:=f(t_{n})$ and $\eta_{n}:=f_{n}(t_{n})^{-1}f_{n}(T_{n})$. By Lemma 2.1, both $\{\gamma_{n}\}$ and $\{\eta_{n}\}$ are escaping sequences in $\Gamma$. So by passing to a subsequence we can suppose that $\gamma_{n}^{-1}\rightarrow x\in\partial_{\infty}X$ and $\eta_{n}\rightarrow y\in\partial_{\infty}X$. Since $t\mapsto f_{n}(t_{n})^{-1}f_{n}(t)$ is a geodesic in $X$ passing through $\operatorname{\mathrm{id}}$ and joining $\gamma_{n}^{-1}$ to $\eta_{n}$, we must have $x\neq y$.
Let $\xi:\partial_{\infty}X\rightarrow\Lambda_{\theta}(\Gamma)$ be the limit map. Proposition 4.3 implies that $U_{\theta}(\gamma_{n}^{-1})\rightarrow\xi(x)$ and $U_{\theta}(\eta_{n})\rightarrow\xi(y)$. Since $\xi(x)\neq\xi(y)$, Proposition 5.1 implies that there exists $C>0$ such that
$\left\|\kappa_{\theta}(f_{n}(T_{n}))-\kappa_{\theta}(f_{n}(t_{n}))-\kappa_{\theta}(f_{n}(t_{n})^{-1}f_{n}(T_{n}))\right\|=\left\|\kappa_{\theta}(\gamma_{n}\eta_{n})-\kappa_{\theta}(\gamma_{n})-\kappa_{\theta}(\eta_{n})\right\|\leq C$
for all $n\geq 1$. So we have a contradiction. ∎
6. Resolution of singularities
In this section we study the asymptotic behavior of proper positive functions which are products of powers of rational functions. This is the key technical step needed to prove that the Poincaré series of a peripheral subgroup of a relatively Anosov subgroup diverges at its critical exponent, see Theorem 7.1.
[…]
The main result of this section is the following theorem. Its proof is motivated by arguments of Benoist–Oh [3, Prop. 7.2], which implies the special case when $m=1$.
Theorem 6.1.
Suppose $R_{1},\dots,R_{m}$ are rational functions on $\operatorname{\mathbb{R}}^{d}$ which are positive and everywhere defined, and let
$R:=R_{1}^{\ell_{1}}\cdots R_{m}^{\ell_{m}}$
where $\ell_{1},\dots,\ell_{m}\in\operatorname{\mathbb{R}}$. If $R$ is a proper function, then:
  1. (1)
    There exists $\delta=\delta(R)>0$ such that
[…]
    converges when $s\in(\delta,+\infty)$ and diverges when $s\in[0,\delta]$.
  2. (2)
    There exist $c=c(R)>0$ and $\epsilon=\epsilon(R)>0$ such that
    $R(x)\geq c(1+\left\|x\right\|)^{\epsilon}$
    for all $x\in\operatorname{\mathbb{R}}^{d}$.
To prove Theorem 6.1 , we compactify $\operatorname{\mathbb{R}}^{d}$ by identifying it with the affine subspace
$\operatorname{\mathbb{A}}_{1}:=\{[x_{0}:\cdots:x_{d}]:x_{0}\neq 0\}\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d+1})$
via the coordinate chart $\psi_{1}:\operatorname{\mathbb{A}}_{1}\to\operatorname{\mathbb{R}}^{d}$ given by
$\psi_{1}:[1:x_{1}:\cdots:x_{d}]\mapsto\left(x_{1},\dots,x_{d}\right).$
It now suffices to prove the following lemma, which is the analog of Theorem 6.1 on “neighborhoods of infinity.”
Lemma 6.2.
For each $p\in\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d+1})-\operatorname{\mathbb{A}}_{1}$, there exist $\delta_{p},c_{p},\epsilon_{p}>0$ and an open neighborhood $\mathcal{O}_{p}\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d+1})$ of $p$ such that:
  1. (1)
    The integral
[…]
    converges when $s\in(\delta_{p},+\infty)$ and diverges when $s\in[0,\delta_{p}]$.
  2. (2)
    $R(x)\geq c_{p}(1+\left\|x\right\|)^{\epsilon_{p}}$ for all $x\in\psi_{1}(\mathcal{O}_{p}\cap\operatorname{\mathbb{A}}_{1})\subset\operatorname{\mathbb{R}}^{d}$.
Assuming Lemma 6.2, we prove Theorem 6.1.
Proof of Theorem 6.1.
Since $\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d+1})-\operatorname{\mathbb{A}}_{1}$ is compact, there exist finitely many points $p_{1},\dots,p_{n}$ in $\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d+1})-\operatorname{\mathbb{A}}_{1}$ such that
$\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d+1})-\operatorname{\mathbb{A}}_{1}\subset\bigcup_{j=1}^{n}\mathcal{O}_{p_{j}}.$
Proof of (1). Let $\delta:=\max\{\delta_{p_{j}}:1\leq j\leq n\}$. Since $R$ is positive,
[…]
Proof of (2). Let $\epsilon:=\min\{\epsilon_{p_{j}}:1\leq j\leq n\}$. Since $R$ is positive and continuous on the compact set $K:=\psi_{1}(\operatorname{\mathbb{A}}_{1}-\bigcup_{j=1}^{n}\mathcal{O}_{p_{j}})\subset\operatorname{\mathbb{R}}^{d}$, there exists $c_{0}>0$ such that
$R(x)\geq c_{0}(1+\left\|x\right\|)^{\epsilon}$
for all $x\in K$. Thus, if we set $c:=\min\left(\{c_{0}\}\cup\{c_{p_{j}}:1\leq j\leq n\}\right)$, then Claim (2) of Lemma 6.2 implies that
$R(x)\geq c(1+\left\|x\right\|)^{\epsilon}$
for all $x\in\operatorname{\mathbb{R}}^{d}$. ∎
To prove Lemma 6.2, first note that we can assume that $p=[0:1:0:\dots:0]$ by changing coordinates. Then $p$ lies in the affine subspace
$\operatorname{\mathbb{A}}_{2}:=\left\{[x_{0}:\cdots:x_{d}]:x_{1}\neq 0\right\}\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d+1}).$
Let $\psi_{2}:\operatorname{\mathbb{A}}_{2}\to\operatorname{\mathbb{R}}^{d}$ be the coordinate chart given by
$\psi_{2}:[y_{1}:1:y_{2}:\dots:y_{d}]\mapsto(y_{1},\dots,y_{d}).$
Observe that $\psi_{2}(p)=0$ and
$\psi_{2}(\operatorname{\mathbb{A}}_{2}-\operatorname{\mathbb{A}}_{1})=\mathcal{Z}:=\left\{(y_{1},\dots,y_{d})\in\operatorname{\mathbb{R}}^{d}:y_{1}=0\right\}.$ (3)
Thus, the restriction
$\psi_{1}\circ\psi_{2}^{-1}|_{\operatorname{\mathbb{R}}^{d}-\mathcal{Z}}:\operatorname{\mathbb{R}}^{d}-\mathcal{Z}\to\operatorname{\mathbb{R}}^{d}$
is a well-defined embedding.
Since $\psi_{1}\circ\psi_{2}^{-1}|_{\operatorname{\mathbb{R}}^{d}-\mathcal{Z}}$ is given by
$\psi_{1}\circ\psi_{2}^{-1}(y_{1},\dots,y_{d})=\left(\frac{1}{y_{1}},\frac{y_{2}}{y_{1}},\dots,\frac{y_{d}}{y_{1}}\right),$ (4)
for all $j\in\{1,\dots,m\}$,
$T_{j}:=R_{j}\circ\psi_{1}\circ\psi_{2}^{-1}:\operatorname{\mathbb{R}}^{d}-\mathcal{Z}\to\operatorname{\mathbb{R}},$
is a rational function that is well-defined and positive on $\operatorname{\mathbb{R}}^{d}-\mathcal{Z}$, and the assumption that $R$ is proper implies that $T:=T_{1}^{\ell_{1}}\dots T_{m}^{\ell_{m}}$ satisfies $\lim_{y\to z}T(y)=\infty$ for all $z\in\mathcal{Z}$. Also, (4) implies that the Jacobian $D(\psi_{1}\circ\psi_{2}^{-1})$ of $\psi_{1}\circ\psi_{2}^{-1}$ satisfies
$\left|\det D(\psi_{1}\circ\psi_{2}^{-1})\right|=\frac{1}{\left|y_{1}\right|^{d+1}},$
and
$\left\|\psi_{1}\circ\psi_{2}^{-1}(y_{1},\ldots,y_{d})\right\|=\sqrt{\frac{1+y_{2}^{2}+\dots+y_{d}^{2}}{y_{1}^{2}}}.$
Thus, to prove Lemma 6.2, it now suffices to prove the following lemma.
Lemma 6.3.
Suppose that $T_{1},\dots,T_{m}$ are rational functions on $\operatorname{\mathbb{R}}^{d}$ which are positive and defined on $\operatorname{\mathbb{R}}^{d}-\mathcal{Z}$, and let
$T:=T_{1}^{\ell_{1}}\dots T_{m}^{\ell_{m}}$
for some $\ell_{1},\dots,\ell_{m}\in\operatorname{\mathbb{R}}$. If $\lim_{y\to z}T(y)=\infty$ for all $z\in\mathcal{Z}$, then there exist $\delta,c,\epsilon>0$ and an open neighborhood $\mathcal{O}\subset\operatorname{\mathbb{R}}^{d}$ of 00 such that:
  1. (1)
    The integral
    $\int_{\mathcal{O}-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda$
    converges when $s\in(\delta,+\infty)$ and diverges when $s\in[0,\delta]$.
  2. (2)
    For all $y=(y_{1},\dots,y_{d})\in\mathcal{O}-\mathcal{Z}$, we have
    $T(y)\geq c\left(1+\sqrt{\frac{1+y_{2}^{2}+\dots+y_{d}^{2}}{y_{1}^{2}}}\right)^{\epsilon}.$
For the remainder of this section, we will focus on the proof of Lemma 6.3. An important tool used in the proof is the following version of Hironaka’s theorem [23] on the resolution of singularities (as stated in [1, pg. 147]).
Theorem 6.4 (Resolution theorem).
Let $F$ be a real analytic function defined in a neighborhood of $0\in\operatorname{\mathbb{R}}^{d}$, and let $\mathcal{Z}$ be the set of zeroes of $F$. If $F$ is not identically zero, then there exists a neighborhood $U$ of 00 in $\operatorname{\mathbb{R}}^{d}$, a real analytic manifold $M$ and a proper real analytic map $\Phi:M\rightarrow U$ such that
  1. (1)
    $\Phi$ restricts to a real-analytic diffeomorphism
    $M-\Phi^{-1}(\mathcal{Z})\rightarrow U-\mathcal{Z}.$
  2. (2)
    For every $q\in M$ there exists an open neighborhood $V_{q}\subset M$ of $q$ and real analytic local coordinates $z_{q,1},\dots,z_{q,d}$ on $V_{q}$ centered at $q$ where
    $F\circ\Phi(z_{q,1},\dots,z_{q,d})=z_{q,1}^{k_{q,1}}\cdots z_{q,d}^{k_{q,d}}\cdot\hat{F}_{q}(z_{q,1},\dots,z_{q,d})$
    for some $k_{q,1},\dots,k_{q,d}\in\operatorname{\mathbb{Z}}_{\geq 0}$ and some nowhere-vanishing real analytic function $\hat{F}_{q}:V_{q}\to\operatorname{\mathbb{R}}$.
The main idea of the proof of Claim (1) of Lemma 6.3 is to try to factor the integrand $\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}$ into the product of a bounded, nowhere vanishing function on $\mathcal{O}$, and power functions on $\mathcal{O}-\mathcal{Z}$ that are “responsible" for how $\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}$ goes to infinity or zero near $\mathcal{Z}$. Doing so allows us to compare the required integral with the integral of a product of power functions, whose convergence or lack thereof is well-understood. Unfortunately, such a global factorization of $\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}$ on $\mathcal{O}$ is not possible in general. However, Theorem 6.4 ensures that we do have such a factorization locally, but at the cost of pre-composing the integrand with a given real analytic function $\Phi$. We will show that the complications introduced to this strategy by using $\Phi$ are surmountable, and that it can indeed be used to prove Lemma 6.3.
We first write, for each $j\in\{1,\dots,n\}$, the rational function $T_{j}$ as
$T_{j}=\frac{f_{j}}{g_{j}}$
where $f_{j},g_{j}:\operatorname{\mathbb{R}}^{d}\to\operatorname{\mathbb{R}}$ are polynomials whose zeroes lie in the set $\mathcal{Z}$ defined by (3). Then
$F:=f_{1}\cdots f_{m}\cdot g_{1}\cdots g_{m}\cdot y_{1}:\operatorname{\mathbb{R}}^{d}\to\operatorname{\mathbb{R}},$ (5)
is a polynomial whose set of zeroes is $\mathcal{Z}$. Applying Theorem 6.4 to this $F$, we get an open neighborhood $U$ of $0\in\mathbb{R}^{d}$, a real analytic manifold $M$, a proper real analytic map $\Phi:M\to U$ such that for all $q\in M$, there exists an open neighborhood $V_{q}$ of $q$ with local coordinates $\{z_{q,r}\}_{r=1}^{d}$ centered at $q$ such that $F\circ\Phi|_{V_{q}}=z_{q,1}^{k_{q,1}}\cdots z_{q,d}^{k_{q,d}}\cdot\hat{F}_{q}$ and $\hat{F}_{q}:V_{q}\to\mathbb{R}$ is a nowhere-vanishing real analytic function. Notice that, via the local coordinates $(z_{q,1},\dots,z_{q,d})$ on $V_{q}$, the Lebesgue measure $\lambda$ on $\operatorname{\mathbb{R}}^{d}$ induces a measure $\lambda_{q}$ on $V_{q}$.
Fix a neighborhood $\mathcal{O}\subset\operatorname{\mathbb{R}}^{d}$ of 00 whose closure is compact and lies in $U$. Also, for each $q\in M$, fix an open neighborhood $V_{q}^{\prime}$ of $q$ whose closure lies in $V_{q}$. Since $\Phi$ is proper, we can find finitely many points $q_{1},\dots,q_{n}\in M$ such that
$\Phi^{-1}(\mathcal{O})\subset\bigcup_{i=1}^{n}V_{q_{i}}^{\prime}.$
For all $i\in\{1,\dots,n\}$ and $r\in\{1,\dots,d\}$, set
$V_{i}:=V_{q_{i}}^{\prime}\cap\Phi^{-1}(\mathcal{O})\subset V_{q_{i}},\quad\hat{F}_{i}:=\hat{F}_{q_{i}}:V_{q_{i}}\rightarrow\operatorname{\mathbb{R}},\quad z_{i,r}:=z_{q_{i},r},\quad k_{i,r}:=k_{q_{i},r}\quad\text{and}\quad\lambda_{i}:=\lambda_{q_{i}}.$
Notice that by construction,
$\overline{V_{i}}\subset V_{q_{i}}$ (6)
is compact, and
$\mathcal{O}=\bigcup_{i=1}^{n}\Phi(V_{i}).$ (7)
Remark 6.5.
In the above set up, one might be tempted to say that by shrinking $\mathcal{O}$, one can find some $q\in M$ such that $\mathcal{O}\subset\Phi(V_{q}^{\prime})$. However, this might not be possible: Each $\Phi(V_{q})\subset U$ might not contain any open neighborhoods of $\Phi(q)$ even if $V_{q}\subset M$ is an open neighborhood of $q$.
Using the sets $V_{i}$ and the map $\Phi$, we have the following local criterion for when the integral
$\int_{\mathcal{O}-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda$
converges.
Lemma 6.6.
For any $s\geq 0$, the integral
$\int_{\mathcal{O}-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda$
converges if and only if the integral
$\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\frac{(T\circ\Phi)^{-s}}{\left|y_{1}\circ\Phi\right|^{d+1}}\left|\det D(\Phi)\right|{\rm d}\lambda_{i}$
converges for all $i\in\{1,\dots,n\}$. Here, $D(\Phi)$ is the Jacobian of $\Phi$ restricted to $V_{q_{i}}$, with respect to the local coordinates $(z_{i,1},\dots,z_{i,d})$ of $V_{q_{i}}$.
Proof.
By (7),
$\max_{i=1,\dots,n}\int_{\Phi(V_{i})-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda\leq\int_{\mathcal{O}-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda\leq\sum_{i=1}^{n}\int_{\Phi(V_{i})-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda.$
Since $\Phi$ restricts to a diffeomorphism $M-\Phi^{-1}(\mathcal{Z})\rightarrow U-\mathcal{Z}$, we have
$\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\frac{(T\circ\Phi)^{-s}}{\left|y_{1}\circ\Phi\right|^{d+1}}\left|\det D(\Phi)\right|{\rm d}\lambda_{i}=\int_{\Phi(V_{i})-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda$
for each $i\in\{1,\dots,n\}$. The lemma follows. ∎
In light of Lemma 6.6, we now need to understand for which values of $s\geq 0$ the integral
$\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\frac{(T\circ\Phi)^{-s}}{\left|y_{1}\circ\Phi\right|^{d+1}}\left|\det D(\Phi)\right|{\rm d}\lambda_{i}$
converges. We do so using the local expressions of the real analytic functions $f_{j}\circ\Phi|_{V_{q_{i}}}$, $g_{j}\circ\Phi|_{V_{q_{i}}}$, $y_{1}\circ\Phi|_{V_{q_{i}}}$ and $\det D(\Phi)|_{V_{q_{i}}}$. By factoring their Taylor series in the local coordinates $(z_{i,1},\dots,z_{i,d})$ on $V_{q_{i}}$, we can write
  • $f_{j}\circ\Phi|_{V_{q_{i}}}=z_{i,1}^{a_{i,j,1}}\cdots z_{i,d}^{a_{i,j,d}}\cdot\hat{f}_{i,j}$ where $a_{i,j,1},\dots,a_{i,j,d}\in\operatorname{\mathbb{Z}}_{\geq 0}$ and $\hat{f}_{i,j}:V_{q_{i}}\to\operatorname{\mathbb{R}}$ is not identically zero on
    $\mathcal{Z}_{i,r}:=\{(z_{i,1},\dots,z_{i,d})\in V_{q_{i}}:z_{i,r}=0\}$
    for any $r\in\{1,\dots,d\}$.
  • $g_{j}\circ\Phi|_{V_{q_{i}}}=z_{i,1}^{b_{i,j,1}}\cdots z_{i,d}^{b_{i,j,d}}\cdot\hat{g}_{i,j}$ where $b_{i,j,1},\dots,b_{i,j,d}\in\operatorname{\mathbb{Z}}_{\geq 0}$ and $\hat{g}_{i,j}:V_{q_{i}}\to\operatorname{\mathbb{R}}$ is not identically zero on $\mathcal{Z}_{i,r}$ for any $r\in\{1,\dots,d\}$.
  • $y_{1}\circ\Phi|_{V_{q_{i}}}=z_{i,1}^{c_{i,1}}\cdots z_{i,d}^{c_{i,d}}\cdot\hat{h}_{i}$ where $c_{i,1},\dots,c_{i,d}\in\operatorname{\mathbb{Z}}_{\geq 0}$ and $\hat{h}_{i}:V_{q_{i}}\to\operatorname{\mathbb{R}}$ is not identically zero on $\mathcal{Z}_{i,r}$ for any $r\in\{1,\dots,d\}$.
  • $\det D(\Phi)|_{V_{q_{i}}}=z_{i,1}^{\gamma_{i,1}}\cdots z_{i,d}^{\gamma_{i,d}}\cdot\hat{J}_{i}$, where $\gamma_{i,1},\dots,\gamma_{i,d}\in\operatorname{\mathbb{Z}}_{\geq 0}$ and $\hat{J}_{i}:V_{q_{i}}\to\operatorname{\mathbb{R}}$ is not identically zero on $\mathcal{Z}_{i,r}$ for any $r\in\{1,\dots,d\}$.
Using Theorem 6.4, we deduce the following lemma about the functions $\hat{f}_{i,j}$, $\hat{g}_{i,j}$, $\hat{h}_{i}$ and
$\hat{W}_{i,s}:=\left(\frac{\hat{g}_{i,1}^{\ell_{1}}\cdots\hat{g}_{i,m}^{\ell_{m}}}{\hat{f}_{i,1}^{\ell_{1}}\cdots\hat{f}_{i,m}^{\ell_{m}}}\right)^{s}\frac{\hat{J}_{i}}{\hat{h}_{i}^{d+1}}$
for all $s\geq 0$.
Lemma 6.7.
Fix $i\in\{1,\dots,n\}$.
  1. (1)
    $\hat{F}_{i}=\hat{f}_{i,1}\cdots\hat{f}_{i,m}\cdot\hat{g}_{i,1}\cdots\hat{g}_{i,m}\cdot\hat{h}_{i}$. In particular, $\hat{f}_{i,1}$, $\dots$, $\hat{f}_{i,m}$, $\hat{g}_{i,1}$, $\dots$, $\hat{g}_{i,m}$ and $\hat{h}_{i}$ are all nowhere-vanishing on $V_{q_{i}}$.
  2. (2)
    For any $s\geq 0$, $\hat{W}_{i,s}$ is a real analytic function which is bounded on $V_{i}$ and not identically zero on $\mathcal{Z}_{i,r}$ for any $r\in\{1,\dots,d\}$.
Proof.
Proof of (1). Observe that
$\frac{\hat{f}_{i,1}\cdots\hat{f}_{i,m}\cdot\hat{g}_{i,1}\cdots\hat{g}_{i,m}\cdot\hat{h}_{i}}{\hat{F}_{i}}=\prod_{r=1}^{d}z_{i,r}^{k_{i,r}-c_{i,r}-\sum_{j=1}^{m}(a_{i,j,r}+b_{i,j,r})}.$
By Theorem 6.4, $\hat{F}_{i}$ is nowhere vanishing, so the left hand side is finite at every point in $V_{q_{i}}$. Thus
$k_{i,r}-c_{i,r}-\sum_{j=1}^{m}(a_{i,j,r}+b_{i,j,r})\geq 0$
for all $r\in\{1,\dots,d\}$. Also, by the definition of $\hat{f}_{i,j}$, $\hat{g}_{i,j}$ and $\hat{h}_{i}$, the left hand side is not identically zero on $\mathcal{Z}_{i,r}$ for all $r\in\{1,\dots,d\}$. Hence we must also have
$k_{i,r}-c_{i,r}-\sum_{j=1}^{m}(a_{i,j,r}+b_{i,j,r})\leq 0$
for all $r\in\{1,\dots,d\}$. So $\hat{F}_{i}=\hat{f}_{i,1}\cdots\hat{f}_{i,m}\cdot\hat{g}_{i,1}\cdots\hat{g}_{i,m}\cdot\hat{h}_{i}$.
Proof of (2). By (1) and the definition of $\hat{J}_{i}$, the function $\hat{W}_{i,s}$ is a real analytic function on $V_{q_{i}}$ which is not identically zero on $\mathcal{Z}_{i,r}$ for any $r\in\{1,\dots,d\}$. Since $\overline{V_{i}}\subset V_{q_{i}}$ is a compact subset (see Equation (6)), the function $\hat{W}_{i,s}$ is bounded on $V_{i}$. ∎
For all $i\in\{1,\dots,n\}$ and $r\in\{1,\dots,d\}$, set
$\beta_{i,r}:=\sum_{j=1}^{m}\ell_{j}(b_{i,j,r}-a_{i,j,r}).$
Since $T$ is positive on $\operatorname{\mathbb{R}}^{d}-\mathcal{Z}$, we can take the absolute value of each term to conclude that
$\frac{(T\circ\Phi)^{-s}}{\left|y_{1}\circ\Phi\right|^{d+1}}\left|\det D(\Phi)\right|$ $=\left|\hat{W}_{i,s}\right|\prod_{r=1}^{d}|z_{i,r}|^{s\beta_{i,r}-(d+1)c_{i,r}+\gamma_{i,r}}.$ (8)
The next lemma characterizes when integrals of functions with the above form converge.
Lemma 6.8.
Suppose $i\in\{1,\dots,n\}$ and $W:V_{i}\rightarrow\operatorname{\mathbb{R}}$ is a bounded real analytic function which is not identically zero on the hyperplane $\mathcal{Z}_{i,r}$ for any $r\in\{1,\dots,d\}$. Then for $\eta_{1},\dots,\eta_{d}\in\operatorname{\mathbb{R}}$ the integral
$\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\left|z_{i,1}\right|^{\eta_{1}}\cdots\left|z_{i,d}\right|^{\eta_{d}}\left|W\right|{\rm d}\lambda_{i}$
converges if and only if $\eta_{r}>-1$ for all $r\in\{1,\dots,d\}$. In particular,
$\int_{\mathcal{O}-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda$
converges if and only if $s\beta_{i,r}-(d+1)c_{i,r}+\gamma_{i,r}>-1$ for all $i\in\{1,\dots,n\}$ and all $r\in\{1,\dots,d\}$.
Proof.
Since $W$ is bounded, it is clear that if $\eta_{r}>-1$ for all $r\in\{1,\dots,d\}$, then the integral converges.
For the other direction, it suffices to assume that $\eta_{1}\leq-1$ and then show that the integral diverges. Since $W$ is not identically zero on $\mathcal{Z}_{i,1}$, we can find an open set $V^{\prime}\subset\mathcal{Z}_{i,1}$ such that $V^{\prime}\subset V_{i}$ and $W$ is nowhere vanishing on $V^{\prime}$. By further shrinking $V^{\prime}$ we can assume that there exists $\epsilon>0$ such that $\overline{(0,\epsilon)\times V^{\prime}}\subset V_{i}$ and
$\overline{V^{\prime}}\cap\bigcup_{r=2}^{d}\mathcal{Z}_{i,r}=\emptyset.$
Then
$\alpha:=\min_{z\in\overline{(0,\epsilon)\times V^{\prime}}}\left|z_{i,2}\right|^{\eta_{2}}\cdots\left|z_{i,d}\right|^{\eta_{d}}\left|W\right|$
is positive. Hence,
$\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}$ $\left|z_{i,1}\right|^{\eta_{1}}\cdots\left|z_{i,d}\right|^{\eta_{d}}\left|W\right|{\rm d}\lambda_{i}\geq\int_{(0,\epsilon)\times V^{\prime}}\left|z_{i,1}\right|^{\eta_{1}}\cdots\left|z_{i,d}\right|^{\eta_{d}}\left|W\right|{\rm d}\lambda_{i}$
[…]
where ${\rm Vol}$ is the measure on $\mathcal{Z}_{i,1}$ induced by the coordinates $(z_{q,2},\dots,z_{q,d})$ on $\mathcal{Z}_{i,1}$ and the Lebesgue measure on $\operatorname{\mathbb{R}}^{d-1}$.
To prove the second statement of the lemma, first note that by Lemma 6.7 Claim (2), $\hat{W}_{i,s}$ is a bounded real analytic function on $V_{i}$ and not identically zero on the hyperplane $\mathcal{Z}_{i,r}$ for any $r\in\{1,\dots,d\}$. So, we may apply the first statement of the lemma to Equation (8) and deduce that
$\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\frac{(T\circ\Phi)^{-s}}{\left|y_{1}\circ\Phi\right|^{d+1}}\left|\det D(\varphi_{i})\right|{\rm d}\lambda_{i}$
converges if and only if
$s\beta_{i,r}-(d+1)c_{i,r}+\gamma_{i,r}>-1$
for all $r\in\{1,\dots,d\}$. The second statement of the lemma now follows from Lemma 6.6. ∎
By Lemma 6.8, we now need to know the values of $s$ so that $s\beta_{i,r}-(d+1)c_{i,r}+\gamma_{i,r}>-1$ for all $i\in\{1,\dots,n\}$ and all $r\in\{1,\dots,d\}$. To that end, it is useful to have the following relations between the exponents $\beta_{i,r}$, $c_{i,r}$ and $\gamma_{i,r}$.
Lemma 6.9.
  1. (1)
    For all $i\in\{1,\dots,n\}$ and all $r\in\{1,\dots,d\}$, $\beta_{i,r}\geq 0$.
  2. (2)
    For all $i\in\{1,\dots,n\}$, $\{r:c_{i,r}>0\}\subset\{r:\beta_{i,r}>0\}$.
  3. (3)
    There exist $i\in\{1,\dots,n\}$ and $r\in\{1,\dots,d\}$ such that
    $c_{i,r}-\gamma_{i,r}\geq 1.$
In particular,
$\delta:=\max\left\{\frac{(d+1)c_{i,r}-\gamma_{i,r}-1}{\beta_{i,r}}:i\in\{1,\dots,n\},\,r\in\{1,\dots,d\}\text{ such that }\beta_{i,r}>0\right\}$
is a positive real number.
Proof.
Proof of (1). Since $T$ is positive on $U-\mathcal{Z}$ and $\lim_{y\to p}T(y)=\infty$ for all $p\in\mathcal{Z}$, the function
$\frac{1}{T\circ\Phi}:M-\Phi^{-1}(\mathcal{Z})\to\operatorname{\mathbb{R}}$
extends to a continuous function on all of $M$. In particular, for all $i\in\{1,\dots,n\}$,
$\frac{1}{T\circ\Phi}\bigg{|}_{V_{i}}=z_{i,1}^{\beta_{i,1}}\cdots z_{i,d}^{\beta_{i,d}}\frac{\hat{g}_{i,1}^{\ell_{1}}\cdots\hat{g}_{i,m}^{\ell_{m}}}{\hat{f}_{i,1}^{\ell_{1}}\cdots\hat{f}_{i,m}^{\ell_{m}}}$ (9)
is finite. By Claim (1) of Lemma 6.7 and the fact that $\overline{V_{i}}\subset V_{q_{i}}$ is compact, we see that $\frac{\hat{g}_{i,1}^{\ell_{1}}\cdots\hat{g}_{i,m}^{\ell_{m}}}{\hat{f}_{i,1}^{\ell_{1}}\cdots\hat{f}_{i,m}^{\ell_{m}}}$ is bounded and nowhere vanishing on $V_{i}$, so (1) follows.
Proof of (2). Since $\lim_{y\to z}T(y)=\infty$ for all $z\in\mathcal{Z}$, it follows that $\frac{1}{T}$ vanishes on $\mathcal{Z}$. Equivalently, $\frac{1}{T\circ\Phi}$ vanishes wherever $y_{1}\circ\Phi$ vanishes. By Lemma 6.7 Claim (1) and the fact that $\overline{V_{i}}\subset V_{q_{i}}$ is compact, $\hat{h}_{i}$ and $\frac{\hat{g}_{i,1}^{\ell_{1}}\cdots\hat{g}_{i,m}^{\ell_{m}}}{\hat{f}_{i,1}^{\ell_{1}}\cdots\hat{f}_{i,m}^{\ell_{m}}}$ are bounded and nowhere vanishing on $V_{i}$, so (2) follows from Equation (9).
Proof of (3). Observe that the integral
$\int_{\mathcal{O}-\mathcal{Z}}\frac{1}{\left|y_{1}\right|}{\rm d}\lambda$
diverges. By (7),
$\int_{\mathcal{O}-\mathcal{Z}}\frac{1}{\left|y_{1}\right|}{\rm d}\lambda\leq\sum_{i=1}^{n}\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\frac{\left|\det D(\Phi)\right|}{|y_{1}\circ\Phi|}{\rm d}\lambda_{i}=\sum_{i=1}^{n}\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\left(\prod_{r=1}^{d}\left|z_{i,r}\right|^{-c_{i,r}+\gamma_{i,r}}\right)\left|\frac{\hat{J}_{i}}{\hat{h}_{i}}\right|{\rm d}\lambda_{i},$
so there is some $i\in\{1,\dots,n\}$ such that the integral
$\int_{V_{i}-\Phi^{-1}(\mathcal{Z})}\left(\prod_{r=1}^{d}\left|z_{i,r}\right|^{-c_{i,r}+\gamma_{i,r}}\right)\left|\frac{\hat{J}_{i}}{\hat{h}_{i}}\right|{\rm d}\lambda_{i}$
diverges. By Lemma 6.7 Claim (1), $\frac{\hat{J}_{i}}{\hat{h}_{i}}$ is real analytic on $V_{q_{i}}$. Also, by the definition of $\hat{J}_{i}$, $\frac{\hat{J}_{i}}{\hat{h}_{i}}$ is not identically zero on $\mathcal{Z}_{i,r}$ for all $r\in\{1,\dots,d\}$. Since $\overline{V_{i}}\subset V_{q_{i}}$ is compact, the function $\frac{\hat{J}_{i}}{\hat{h}_{i}}$ is bounded on $V_{i}$. So by Lemma 6.8 there is some $r\in\{1,\dots,d\}$ such that $-c_{i,r}+\gamma_{i,r}\leq-1$. This proves (3).
We will now deduce the final claim of the lemma. By (3), there is some $i\in\{1,\dots,n\}$ and $r\in\{1,\dots,d\}$ such that $c_{i,r}-\gamma_{i,r}\geq 1$. Since $\gamma_{i,r}\geq 0$, it follows that $c_{i,r}>0$, so (2) implies that $\beta_{i,r}>0$. Then
$\frac{(d+1)c_{i,r}-\gamma_{i,r}-1}{\beta_{i,r}}>0,$
which implies that $\delta>0$. ∎
Combining Lemmas 6.8 and 6.9, we may now prove Lemma 6.3.
Proof of Lemma 6.3.
Proof of (1). Let $\delta>0$ be the quantity specified in the statement of Lemma 6.9.
Suppose that $s\in(\delta,+\infty)$. Pick any $i\in\{1,\dots,n\}$ and $r\in\{1,\dots,d\}$. If $\beta_{i,r}=0$, then Lemma 6.9 Claim (2) implies that $c_{i,r}=0$, in which case
$s\beta_{i,r}-(d+1)c_{i,r}+\gamma_{i,r}=\gamma_{i,r}\geq 0>-1.$
If $\beta_{i,r}\neq 0$, then $s>\delta\geq\frac{(d+1)c_{i,r}-\gamma_{i,r}-1}{\beta_{i,r}}$. By Lemma 6.9 Claim (1), $\beta_{i,r}>0$, so
$s\beta_{i,r}-(d+1)c_{i,r}+\gamma_{i,r}>-1.$
It now follows from Lemma 6.8 that
$\int_{\mathcal{O}-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda$
converges.
Next, suppose that $s\in[0,\delta]$. Then by definition, there is some $i\in\{1,\dots,n\}$ and $r\in\{1,\dots,d\}$ such that
$s\beta_{i,r}-(d+1)c_{i,r}+\gamma_{i,r}\leq-1.$
Then by Lemma 6.8,
$\int_{\mathcal{O}-\mathcal{Z}}\frac{T^{-s}}{\left|y_{1}\right|^{d+1}}{\rm d}\lambda$
diverges. This completes the proof of (1).
Proof of (2). Recall that
$\frac{1}{T\circ\Phi}\bigg{|}_{V_{i}}=z_{i,1}^{\beta_{i,1}}\cdots z_{i,d}^{\beta_{i,d}}R$
where $R=\frac{\hat{g}_{i,1}^{\ell_{1}}\cdots\hat{g}_{i,m}^{\ell_{m}}}{\hat{f}_{i,1}^{\ell_{1}}\cdots\hat{f}_{i,m}^{\ell_{m}}}$ is a nowhere vanishing analytic function (by Lemma 6.7) and that
$y_{1}\circ\Phi|_{V_{q_{i}}}=z_{i,1}^{c_{i,1}}\cdots z_{i,d}^{c_{i,d}}\cdot\hat{h}_{i}$
where $\hat{h}_{i}$ is a nowhere vanishing analytic function (again by Lemma 6.7). Since $\overline{V_{i}}\subset V_{q_{i}}$ is compact, $\beta_{i,r}\geq 0$ for all $r$ and $\beta_{i,r}=0$ whenever $c_{i,r}=0$ (by Lemma 6.9), there exist $C_{i},\epsilon_{i}>0$ such that
$\frac{1}{T\circ\Phi(z)}\leq C_{i}\left|y_{1}\circ\Phi(z)\right|^{\epsilon_{i}}$ (10)
for all $z\in V_{i}$. Set
$\epsilon:=\max\{\epsilon_{1},\dots,\epsilon_{n}\}.$
Since the closure of $\mathcal{O}$ is compact, there exists $c_{0}>0$ such that
$1\geq c_{0}\left(\left|y_{1}\right|+\sqrt{1+y_{2}^{2}+\dots+y_{d}^{2}}\right)$
for all $y\in\mathcal{O}$. Hence
$\frac{1}{\left|y_{1}\right|}\geq c_{0}\left(1+\sqrt{\frac{1+y_{2}^{2}+\dots+y_{d}^{2}}{y_{1}^{2}}}\right)$ (11)
for all $y\in\mathcal{O}-\mathcal{Z}$. Set
$c:=\min\left\{\frac{c_{0}^{\epsilon_{1}}}{C_{1}},\dots,\frac{c_{0}^{\epsilon_{n}}}{C_{n}}\right\}.$
Fix $y\in\mathcal{O}-\mathcal{Z}$. Then there exist $i\in\{1,\dots,m\}$ and $z\in V_{i}$ such that $y=\Phi(z)$. Then
$T(y)\geq\frac{1}{C_{i}\left|y_{1}\right|^{\epsilon_{i}}}\geq c\left(1+\sqrt{\frac{1+y_{2}^{2}+\dots+y_{d}^{2}}{y_{1}^{2}}}\right)^{\epsilon},$
where the first inequality holds by Equation (10) and the second inequality holds by Equation (11).
7. Entropy gap for peripheral subgroups
In this section we prove that the Poincaré series associated to any peripheral subgroup diverges at its critical exponent.
Theorem 7.1 (Theorem 1.11).
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in\mathfrak{a}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$. If $P\in\mathcal{P}$, then $Q_{P}^{\phi}$ diverges at its critical exponent.
Delaying the proof of Theorem 7.1 for a moment, we observe that it implies that the critical exponent of the peripheral subgroup is strictly smaller than the critical exponent of the entire group.
Corollary 7.2.
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in\mathfrak{a}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$. If $P\in\mathcal{P}$, then $\delta^{\phi}(P)<\delta^{\phi}(\Gamma)$.
Proof.
Notice that $\Lambda_{\theta}(P)$ consists of a single point, namely the fixed point of $P$ in $\Lambda_{\theta}(\Gamma)$. Hence Theorems 1.6 and 7.1 imply that $\delta^{\phi}(P)<\delta^{\phi}(\Gamma)$. ∎
The rest of the section is devoted to the proof of Theorem 7.1, so fix $\Gamma$, $\mathcal{P}$ and $\phi$ as in the statement of the theorem.
Let $p:\mathsf{G}\rightarrow\mathsf{G}^{\prime}$ and $\theta^{\prime}\subset\Delta^{\prime}$ be as in Proposition 4.5. By part (4) of that proposition, $\Gamma^{\prime}:=p(\Gamma)$ is a $\operatorname{\mathsf{P}}^{\prime}_{\theta^{\prime}}$-Anosov subgroup relative to $\mathcal{P}^{\prime}:=\{p(P):P\in\mathcal{P}\}$. Also let $\phi^{\prime}\in({\mathfrak{a}}_{\theta^{\prime}}^{\prime})^{*}$ be the unique functional where $\phi^{\prime}\circ{\rm d}p=\phi$. Then $Q_{p(P)}^{\phi^{\prime}}=Q_{P}^{\phi}$ for all $P\in\mathcal{P}$. So by replacing $\mathsf{G}$ with $\mathsf{G}^{\prime}$, we may assume that $\mathsf{G}$ has trivial center, and that $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$.
Fix $P\in\mathcal{P}$. Then by Theorem 4.4 there exists a closed subgroup $\mathsf{H}\subset\mathsf{G}$ with finitely many components such that:
  1. (1)
    $P$ is a cocompact lattice in $\mathsf{H}$.
  2. (2)
    $\mathsf{H}=\mathsf{L}\ltimes\mathsf{U}$ where $\mathsf{L}$ is compact and $\mathsf{U}$ is the unipotent radical of $\mathsf{H}$.
  3. (3)
    $\mathsf{H}^{0}=\mathsf{L}^{0}\times\mathsf{U}$ and $\mathsf{L}^{0}$ is Abelian.
Let $\mathfrak{u}$ denote the Lie algebra of $\mathsf{U}$.
Since $\omega_{\alpha}(\kappa_{\theta}(g))=\omega_{\alpha}(\kappa(g))$ for all $\alpha\in\theta$ and $g\in\mathsf{G}$, by Proposition 2.3, for any $\alpha\in\theta$ there exist $M_{\alpha}\in\operatorname{\mathbb{N}}$, $C_{\alpha}>1$ and a positive everywhere defined rational function $R_{\alpha}:\mathfrak{u}\rightarrow\operatorname{\mathbb{R}}$ where
$\frac{1}{C_{\alpha}}R_{\alpha}(Y)^{1/M_{\alpha}}\leq e^{\omega_{\alpha}(\kappa_{\theta}(\exp(Y)))}\leq C_{\alpha}R_{\alpha}(Y)^{1/M_{\alpha}}$
for all $Y\in\mathfrak{u}$. Write $\phi=\sum_{\alpha\in\theta}c_{\alpha}\omega_{\alpha}$. Then define $R:=\prod_{\alpha\in\theta}R_{\alpha}^{\left|c_{\alpha}\right|/M_{\alpha}}$ and $C_{\phi}:=\prod_{\alpha\in\theta}C_{\alpha}^{\left|c_{\alpha}\right|}$. Note that
$C_{\phi}^{-s}R^{-s}(Y)\leq e^{-s\phi(\kappa_{\theta}(\exp(Y)))}\leq C_{\phi}^{s}R^{-s}(Y)$
for all $s\in\operatorname{\mathbb{R}}$.
Lemma 7.3.
$R:\mathfrak{u}\rightarrow\operatorname{\mathbb{R}}$ is proper.
Proof.
Suppose that $\{Y_{n}\}$ is an escaping sequence in $\mathfrak{u}$. Since $\exp:\mathfrak{u}\rightarrow\mathsf{U}$ is a diffeomorphism, $\{\exp(Y_{n})\}$ is an escaping sequence in $\mathsf{U}$ (see Proposition 2.3). Since $P$ is a cocompact lattice in $\mathsf{H}$, there exists an escaping sequence $\{\gamma_{n}\}$ in $P$ such that $\{\gamma_{n}^{-1}\exp(Y_{n})\}$ is relatively compact in $\mathsf{H}$. So there exists $C_{0}>0$ such that
$\left\|\kappa\left(\gamma_{n}^{-1}\exp(Y_{n})\right)\right\|\leq C_{0}$
for all $n\geq 1$. Then by Lemma 2.1
$R(Y_{n})\geq\frac{1}{C_{\phi}}e^{\phi(\kappa_{\theta}(\exp(Y_{n})))}\geq\frac{1}{C_{\phi}}e^{\phi(\kappa_{\theta}(\gamma_{n}))}e^{-\left\|\phi\right\|C_{0}},$
where $\left\|\phi\right\|$ is the operator norm of the linear map $\phi:\mathfrak{a}\to\operatorname{\mathbb{R}}$. Since $\delta^{\phi}(\Gamma)<+\infty$ and $\{\gamma_{n}\}$ is an escaping sequence, we must have $\phi(\kappa_{\theta}(\gamma_{n}))\rightarrow+\infty$. Hence $R(Y_{n})\rightarrow+\infty$. So $R$ is proper. ∎
Fix a measure $\lambda$ on $\mathfrak{u}$ which is obtained by pulling back the Lebesgue measure on $\operatorname{\mathbb{R}}^{\dim\mathfrak{u}}$ via some linear isomorphism (notice that any two measures obtained this way are scalar multiplies of each other). Then the push-forward $\mu:=\exp_{*}\lambda$ is a Haar measure on $\mathsf{U}$, see Proposition 2.3. By Theorem 6.1 there exists $\delta>0$ such that
[…]
converges when $s\in(\delta,+\infty)$ and diverges when $s\in[0,\delta]$. Hence to complete the proof of Theorem 7.1 it suffices to show the following.
Lemma 7.4.
There exists a continuous function $A:\operatorname{\mathbb{R}}_{\geq 0}\to\operatorname{\mathbb{R}}_{>0}$ such that
$\frac{1}{A(s)}Q_{P}^{\phi}(s)\leq Q_{R}(s)\leq A(s)Q_{P}^{\phi}(s)$
for all $s\geq 0$.
Proof.
We prove the lemma via a series of estimates. First let
$P_{0}:=\mathsf{H}^{0}\cap P=\left(\mathsf{L}^{0}\times\mathsf{U}\right)\cap P.$
Since $\mathsf{H}$ has finitely many connected components, $P_{0}$ has finite index in $P$. Let $\gamma_{1},\dots,\gamma_{n}\in P$ such that $P/P_{0}=\{\gamma_{1}P_{0},\dots,\gamma_{n}P_{0}\}$, and let
$D:=\left\|\phi\right\|\max_{i=1,\dots,n}\left\|\kappa(\gamma_{i})\right\|,$
where $\left\|\phi\right\|$ is the operator norm of the linear map $\phi:\mathfrak{a}_{\theta}\to\operatorname{\mathbb{R}}$. Then by Lemma 2.1,
$\frac{e^{-Ds}}{n}Q_{P}^{\phi}(s)\leq Q_{P_{0}}^{\phi}(s)\leq Q_{P}^{\phi}(s)$ (12)
for all $s\geq 0$.
Next, let $\pi:\mathsf{L}^{0}\times\mathsf{U}\rightarrow\mathsf{U}$ denote the projection and let $P_{1}:=\pi(P_{0})$. Since $P_{0}$ is discrete and $\mathsf{L}^{0}$ is compact, the kernel of $\pi|_{P_{0}}$ is finite and $P_{1}$ is discrete. Then by Lemma 2.1,
$\frac{e^{-Cs}}{B}Q_{P_{0}}^{\phi}(s)\leq Q_{P_{1}}^{\phi}(s)\leq Q_{P_{0}}^{\phi}(s)$ (13)
for all $s\geq 0$, where $B:=\#\ker(\pi|_{P_{0}})$ and
$C:=\left\|\phi\right\|\max_{g\in\mathsf{L}^{0}}\left\|\kappa(g)\right\|.$
By (12) and (13), it now suffices to show that there exists a continuous function $A:\operatorname{\mathbb{R}}_{\geq 0}\to\operatorname{\mathbb{R}}_{>0}$ such that
$\frac{1}{A(s)}Q_{P_{1}}^{\phi}(s)\leq Q_{R}(s)\leq A(s)Q_{P_{1}}^{\phi}(s)$ (14)
for all $s\geq 0$.
Since $P$ is a cocompact lattice in $\mathsf{H}$, we see that $P_{1}$ is a cocompact lattice in $\mathsf{U}$. Fix compact neighborhoods $K_{1},K_{2}\subset\mathsf{U}$ of the identity such that the left $P_{1}$-translates of $K_{1}$ are pairwise disjoint and the left $P_{1}$-translates of $K_{2}$ cover $\mathsf{U}$. Since $K_{1}$ and $K_{2}$ have non-empty interior, both $\mu(K_{1})$ and $\mu(K_{2})$ are positive. For both $i=1,2$, define
$E_{i}:=\left\|\phi\right\|\max_{h\in K_{i}}\left\|\kappa(h)\right\|.$
Then by Lemma 2.1,
$e^{-sE_{i}}e^{-s\phi(\kappa_{\theta}(g))}\leq e^{-s\phi(\kappa_{\theta}(\gamma))}\leq e^{sE_{i}}e^{-s\phi(\kappa_{\theta}(g))}$
for all $\gamma\in P_{1}$, $g\in\gamma K_{i}$ and $s\geq 0$. So
[…]
for all $\gamma\in P_{1}$ and $s>0$. Then
[…]
Likewise,
$Q_{R}(s)$ $\geq C_{\phi}^{-s}\mu(K_{1})e^{-sE_{1}}Q_{P_{1}}^{\phi}(s),$
so (14) holds. ∎
Remark 7.5.
Notice that the proof implies that $\delta^{\phi}(P)$ depends only on $\phi$ and $\mathsf{U}$. One may concisely describe $\mathsf{U}$ as the unipotent radical of the Zariski closure of $P$.
8. Divergence of the Poincaré series at its critical exponent
In this section we use an argument of Dal’bo–Otal–Peigné [17] and Corollary 7.2 to prove that the Poincaré series diverges at its critical exponent. A similar argument was used by Blayac–Zhu [5] in their study of relatively Anosov subgroups of the projective general linear group which preserve a properly convex domain and Patterson–Sullivan measures defined using the Busemann functions associated to the Hilbert distance on the properly convex domain.
Theorem 8.1 (Theorem 1.1).
If $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in\mathfrak{a}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$, then $Q_{\Gamma}^{\phi}$ diverges at its critical exponent.
The key result needed to prove Theorem 8.1 is the following lemma.
Lemma 8.2.
If $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in\mathfrak{a}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$, then there exists a $\phi$-Patterson–Sullivan measure $\mu$ for $\Gamma$ of dimension $\delta^{\phi}(\Gamma)$ such that $\mu$ has no atoms.
Assuming Lemma 8.2, we prove Theorem 8.1.
Proof of Theorem 8.1 assuming Lemma 8.2.
By Lemma 8.2 there exists an atom-less $\phi$-Patterson–Sullivan measure $\mu$ of dimension $\delta^{\phi}(\Gamma)$. Suppose for a contradiction that $Q_{\Gamma}^{\phi}(\delta^{\phi}(\Gamma))<+\infty$. Then Theorem 1.4 implies that $\mu(\Lambda_{\theta}^{\rm con}(\Gamma))=0$. Since $\Lambda_{\theta}(\Gamma)-\Lambda_{\theta}^{\rm con}(\Gamma)$ is countable and $\mu$ has no atoms, $\mu(\Lambda_{\theta}(\Gamma))=0$. Since $\mu$ is a probability measure on $\Lambda_{\theta}(\Gamma)$ this is impossible. ∎
We now give the proof of Lemma 8.2.
Proof of Lemma 8.2.
Let $\delta:=\delta^{\phi}(\Gamma)$. Endow $\Gamma\cup\Lambda_{\theta}(\Gamma)$ with the topology from Lemma 4.1 and for $x\in\Gamma\cup\Lambda_{\theta}(\Gamma)$ let $\mathcal{D}_{x}$ denote the Dirac measure supported on $x$. By [36, Lem. 3.1] there exists a continuous non-decreasing function $h:\operatorname{\mathbb{R}}^{+}\to\operatorname{\mathbb{R}}^{+}$ such that:
  1. (1)
    The series
    $\hat{Q}(s):=\sum_{\gamma\in\Gamma}h\left(e^{\phi(\kappa_{\theta}(\gamma))}\right)e^{-s\phi(\kappa_{\theta}(\gamma))}$
    converges for $s>\delta$ and diverges for $s\leq\delta$.
  2. (2)
    For any $\epsilon>0$ there exists $\nu_{0}>0$ such that: if $s>1$ and $\nu>\nu_{0}$, then $h(\nu s)\leq s^{\epsilon}h(\nu)$.
(In the case when $Q_{\Gamma}^{\phi}$ diverges at its critical exponent, we can choose $h\equiv 1$.)
For $s>\delta$ consider the probability measure
$\mu_{s}:=\frac{1}{\hat{Q}(s)}\sum_{\gamma\in\Gamma}h\left(e^{\phi(\kappa_{\theta}(\gamma))}\right)e^{-s\phi(\kappa_{\theta}(\gamma))}\mathcal{D}_{\gamma}$
on $\Gamma\cup\Lambda_{\theta}(\Gamma)$. By compactness, there exists $\{s_{n}\}\subset(\delta,\infty)$ so that $\lim_{n\rightarrow\infty}s_{n}=\delta$ and
$\mu:=\lim_{n\rightarrow\infty}\mu_{s_{n}}$
exists in the weak-*. topology. In the proof of [15, Prop. 3.1], we showed that $\mu$ is a $\phi$-Patterson–Sullivan measure for $\Gamma$ of dimension $\delta$.
We will now prove that $\mu$ has no atoms. By [15, Prop. 8.1], if $F\in\Lambda_{\theta}^{\rm con}(\Gamma)$, then $\mu(\{F\})=0$. So it suffices to fix a parabolic point $F_{0}\in\Lambda_{\theta}(\Gamma)$ and show that $\mu(\{F_{0}\})=0$. By the definition of the weak-* topology, it suffices to find a sequence of open sets $\{V_{m}\}$ in $\Gamma\cup\Lambda_{\theta}(\Gamma)$ each of which contains $F_{0}$, and where
$\liminf_{m\rightarrow\infty}\,\limsup_{s\searrow\delta}\,\mu_{s}(V_{m})=0.$ (15)
Let $\xi:\partial(\Gamma,\mathcal{P})\to\operatorname{\mathcal{F}}_{\theta}$ be the limit map of $\Gamma$. Let $P$ be the stabilizer in $\Gamma$ of $F_{0}$ (it is a conjugate of a subgroup in $\mathcal{P}$), fix an enumeration $P=\{g_{1},g_{2},\dots\}$, and let $Q_{m}:=\{g_{1},\dots,g_{m}\}\subset P$ for each positive integer $m$.
Fix a Groves–Manning cusp space $X$ for $(\Gamma,\mathcal{P})$, and let $\Gamma^{\prime}$ be the set of elements $\gamma\in\Gamma$ for which there is a geodesic ray $\sigma:[0,\infty)\rightarrow X$ such that $\sigma(0)=\gamma$, $\sigma(T)=\operatorname{\mathrm{id}}$ for some $T\geq 0$ and $\lim_{t\rightarrow\infty}\sigma(t)=\xi^{-1}(F_{0})\in\partial_{\infty}X=\partial(\Gamma,\mathcal{P})$. The next lemma implies that for all positive integers $m$,
$V_{m}:=(\Gamma\cup\Lambda_{\theta}(\Gamma))-Q_{m}\cdot\overline{\Gamma^{\prime}}$
is an open neighborhood of $F_{0}$, where $\overline{\Gamma^{\prime}}$ denotes the closure of $\Gamma^{\prime}$ in $\Gamma\cup\Lambda_{\theta}(\Gamma)$.
Lemma 8.3.
The closed subset $\overline{\Gamma^{\prime}}\subset\Gamma\cup\Lambda_{\theta}(\Gamma)$ does not contain $F_{0}$. In particular, for all positive integers $m$, $V_{m}\subset\Gamma\cup\Lambda_{\theta}(\Gamma)$ is an open neighborhood of $F_{0}$.
Proof.
Suppose for a contradiction that there exists a sequence $\{\gamma_{n}\}$ in $\Gamma^{\prime}$ converging to $F_{0}$. Then Lemma 4.1 implies that $U_{\theta}(\gamma_{n})\rightarrow F_{0}$. Hence, if we view $\{\gamma_{n}\}$ as a sequence in $X$, then by Proposition 4.3 and taking a subsequence if necessary, we must have $\gamma_{n}=\gamma_{n}(\operatorname{\mathrm{id}})\rightarrow\xi^{-1}(F_{0})$.
For each $n$ fix a geodesic ray $\sigma_{n}:[0,\infty)\rightarrow X$ such that $\sigma_{n}(0)=\gamma_{n}$, $\sigma_{n}(T_{n})=\operatorname{\mathrm{id}}$ for some $T_{n}\geq 0$ and $\lim_{t\rightarrow\infty}\sigma_{n}(t)=\xi^{-1}(F_{0})$. Passing to a subsequence we can suppose that the geodesics $\{\sigma_{n}(\cdot+T_{n})\}$ converges locally uniformly to a geodesic $\sigma:\operatorname{\mathbb{R}}\rightarrow X$. Then, since $\gamma_{n}\rightarrow\xi^{-1}(F_{0})$, we have
$\lim_{t\rightarrow-\infty}\sigma(t)=\lim_{n\rightarrow\infty}\sigma_{n}(0)=\lim_{n\rightarrow\infty}\gamma_{n}=\xi^{-1}(F_{0})=\lim_{t\rightarrow+\infty}\sigma(t),$
which is impossible.
Since $Q_{m}$ is finite for all $m$, the second claim follows immediately from the first. ∎
Since $P$ is conjugate to a subgroup in $\mathcal{P}$, by Corollary 7.2,
$\epsilon:=\frac{\delta^{\phi}(\Gamma)-\delta^{\phi}(P)}{2}>0.$
To prove that Equation (15) holds, we prove the following lemma. This will be used to give an upper bound on $\mu_{s}(V_{m})$ involving the tail of the Poincaré series $Q_{P}^{\phi}(\delta^{\phi}(P)+\epsilon)$.
Lemma 8.4.
  1. (1)
    $P\cdot\Gamma^{\prime}=\Gamma$.
  2. (2)
    There exists $C_{1}>0$ such that
    $\left|\phi(\kappa_{\theta}(g\gamma))-\phi(\kappa_{\theta}(g))-\phi(\kappa_{\theta}(\gamma))\right|\leq C_{1}$
    for all $\gamma\in\Gamma^{\prime}$ and $g\in P$.
  3. (3)
    There exists $C_{2}>0$ such that
    $h\left(e^{\phi(\kappa_{\theta}(g_{k}))+\phi(\kappa_{\theta}(\gamma))+C_{1}}\right)\leq C_{2}e^{\epsilon\phi(\kappa_{\theta}(g_{k}))}h\left(e^{\phi(\kappa_{\theta}(\gamma))}\right)$
    for all $\gamma\in\Gamma^{\prime}$ and $k\geq 1$.
Proof.
Proof of (1). Fix $\gamma\in\Gamma$. Then fix a geodesic ray $\sigma:[0,\infty)\rightarrow X$ with $\sigma(0)=\gamma$ and $\lim_{t\rightarrow\infty}\sigma(t)=\xi^{-1}(F_{0})$. Let $\mathcal{H}_{P}\subset X$ denote the combinatorial horoball associated to $P$. Notice that the geodesic ray $\sigma_{0}:[0,\infty)\rightarrow X$ which satisfies $\sigma_{0}(n)=(\operatorname{\mathrm{id}},n)\in\mathcal{H}_{P}$ for all $n\in\operatorname{\mathbb{N}}$ also limits to $\xi^{-1}(F_{0})$. Hence
$\sup_{t\geq 0}\,{\rm d}_{X}(\sigma(t),\sigma_{0}(t))<+\infty.$
Since
$\lim_{t\rightarrow\infty}{\rm d}_{X}(\sigma_{0}(t),X-\mathcal{H}_{P})=+\infty,$
there exists $T\geq 0$ such that $\sigma(T)\in P$ and $\sigma(t)\in\mathcal{H}_{P}$ for all $t\geq T$. Then $\sigma(T)^{-1}\gamma\in\Gamma^{\prime}$. Since $\gamma\in\Gamma$ was arbitrary, (1) holds.
Proof of (2). Suppose not. Then for each $n\geq 1$ there exist $\gamma_{n}\in\Gamma^{\prime}$ and $g_{n}\in P$ such that
$\left|\phi(\kappa_{\theta}(g_{n}\gamma_{n}))-\phi(\kappa_{\theta}(g_{n}))-\phi(\kappa_{\theta}(\gamma_{n}))\right|\geq n.$
Lemma 2.1 implies that $\{\gamma_{n}\}$ and $\{g_{n}\}$ are both escaping sequences in $\Gamma$. Since $\{g_{n}\}\subset P$, in the topology on $\Gamma\cup\Lambda_{\theta}(\Gamma)$ we have $g_{n}^{-1}\rightarrow F_{0}$. Passing to a subsequence we can also assume that
$F:=\lim_{n\rightarrow\infty}\gamma_{n}\in\Lambda_{\theta}(\Gamma)$
exists. By Lemma 8.3, $F_{0}\neq F$. So by Lemma 4.1 and Proposition 5.1, we have
$\limsup_{n\rightarrow\infty}\left|\phi(\kappa_{\theta}(g_{n}\gamma_{n}))-\Big{(}\phi(\kappa_{\theta}(g_{n}))+\phi(\kappa_{\theta}(\gamma_{n}))\Big{)}\right|<+\infty$
and hence a contradiction.
Proof of (3). Let $\nu_{0}>0$ be a constant such that if $s>1$ and $\nu>\nu_{0}$, then $h(\nu s)\leq s^{\epsilon}h(\nu)$. Let $C_{1}>0$ be the constant from (2), and fix $C_{1}^{\prime}\geq C_{1}$ such that
$\phi(\kappa_{\theta}(g_{k}))+C_{1}^{\prime}>0\quad\text{and}\quad\phi(\kappa_{\theta}(\gamma))+C_{1}^{\prime}>\log\nu_{0}$
for all $\gamma\in\Gamma^{\prime}$ and $k\geq 1$. Also let
$\Gamma^{\prime\prime}:=\{\gamma\in\Gamma^{\prime}:\phi(\kappa_{\theta}(\gamma))\leq\log\nu_{0}\}.$
If $\gamma\notin\Gamma^{\prime\prime}$, then
$h\left(e^{\phi(\kappa_{\theta}(g_{k}))+\phi(\kappa_{\theta}(\gamma))+C_{1}}\right)\leq h\left(e^{\phi(\kappa_{\theta}(g_{k}))+\phi(\kappa_{\theta}(\gamma))+C_{1}^{\prime}}\right)\leq e^{\epsilon C_{1}^{\prime}}e^{\epsilon\phi(\kappa_{\theta}(g_{k}))}h\left(e^{\phi(\kappa_{\theta}(\gamma))}\right).$
On the other hand, if $\gamma\in\Gamma^{\prime\prime}$, then
$h\left(e^{\phi(\kappa_{\theta}(g_{k}))+\phi(\kappa_{\theta}(\gamma))+C_{1}}\right)\leq h\left(e^{\phi(\kappa_{\theta}(g_{k}))+\phi(\kappa_{\theta}(\gamma))+2C_{1}^{\prime}}\right)\leq e^{\epsilon C_{1}^{\prime}}e^{\epsilon\phi(\kappa_{\theta}(g_{k}))}h\left(e^{\phi(\kappa_{\theta}(\gamma))+C_{1}^{\prime}}\right).$
So (3) holds with
$C_{2}:=e^{\epsilon C_{1}^{\prime}}\max\left\{\frac{h\left(e^{\phi(\kappa_{\theta}(\gamma))+C_{1}^{\prime}}\right)}{h\left(e^{\phi(\kappa_{\theta}(\gamma))}\right)}:\gamma\in\Gamma^{\prime\prime}\right\},$
which is finite since $\Gamma^{\prime\prime}$ is finite. ∎
If $s>\delta$, then by Lemma 8.4,
$\mu_{s}(V_{m})$ $\leq\frac{1}{\hat{Q}(s)}\sum_{k>m}\sum_{\gamma\in\Gamma^{\prime}}h\left(e^{\phi(\kappa_{\theta}(g_{k}\gamma))}\right)e^{-s\phi(\kappa_{\theta}(g_{k}\gamma))}$
$\leq\frac{e^{sC_{1}}}{\hat{Q}(s)}\sum_{k>m}\sum_{\gamma\in\Gamma^{\prime}}h\left(e^{\phi(\kappa_{\theta}(g_{k}))+\phi(\kappa_{\theta}(\gamma))+C_{1}}\right)e^{-s\phi(\kappa_{\theta}(g_{k}))-s\phi(\kappa_{\theta}(\gamma))}$
$\leq\frac{C_{2}e^{sC_{1}}}{\hat{Q}(s)}\sum_{k>m}\sum_{\gamma\in\Gamma^{\prime}}e^{-(s-\epsilon)\phi(\kappa_{\theta}(g_{k}))}h\left(e^{\phi(\kappa_{\theta}(\gamma))}\right)e^{-s\phi(\kappa_{\theta}(\gamma))}$
$\leq C_{2}e^{sC_{1}}\sum_{k>m}e^{-(\delta^{\phi}(P)+\epsilon)\phi(\kappa_{\theta}(g_{k}))}.$
Since $Q_{P}^{\phi}(\delta^{\phi}(P)+\epsilon)<+\infty$, it follows that
$\liminf_{m\rightarrow\infty}\,\limsup_{s\searrow\delta}\,\mu_{s}(V_{m})\leq C_{2}e^{\delta C_{1}}\liminf_{m\rightarrow\infty}\sum_{k>m}e^{-(\delta^{\phi}(P)+\epsilon)\phi(\kappa_{\theta}(g_{k}))}=0,$
so Equation (15) holds. ∎
Remark 8.5 (The elementary case).
In this remark we sketch why the Poincaré series of an infinite elementary transverse groups diverges at its critical exponent.
Suppose $\Gamma\subset\mathsf{G}$ is an infinite $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup, #$\#\Lambda_{\theta}(\Gamma)\leq 2$, $\phi\in\mathfrak{a}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$.
Case 1: Suppose $\Lambda_{\theta}(\Gamma)=\{F^{+},F^{-}\}$. Then one can show that there is an infinite order element $\gamma\in\Gamma$ such that $\langle\gamma\rangle$ has finite index in $\Gamma$. Moreover, we can assume that $\gamma^{n}(F)\rightarrow F^{+}$ for all $F\in\operatorname{\mathcal{F}}_{\theta}$ transverse to $F^{-}$. Then arguing as in Case 2 of the proof of Corollary 9.1 below, one can show that $\delta^{\phi}(\Gamma)=0$. Hence, since $\Gamma$ is infinite, we have
$Q_{\Gamma}^{\phi}(0)=\sum_{\gamma\in\Gamma}1=+\infty.$
Case 2: Suppose $\Lambda_{\theta}(\Gamma)=\{F^{+}\}$. Using Proposition 4.5 we may assume that $\mathsf{G}$ has trivial center, and that $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$. Then one can adapt the proof of Theorem 4.4 part (2) to show that there exists a closed subgroup $\mathsf{H}\subset\mathsf{G}$ with finitely many components such that:
  1. (1)
    $\Gamma$ is a cocompact lattice in $\mathsf{H}$.
  2. (2)
    $\mathsf{H}=\mathsf{L}\ltimes\mathsf{U}$ where $\mathsf{L}$ is compact and $\mathsf{U}$ is the unipotent radical of $\mathsf{H}$.
  3. (3)
    $\mathsf{H}^{0}=\mathsf{L}^{0}\times\mathsf{U}$ and $\mathsf{L}^{0}$ is Abelian.
Finally, one can use the proof of Theorem 7.1 to show that $Q_{\Gamma}^{\phi}$ diverges at its critical exponent.
9. Relatively quasiconvex subgroups
Suppose that $\Gamma$ is hyperbolic relative to $\mathcal{P}$. Given a subgroup $\Gamma_{0}\subset\Gamma$, the limit set of $\Gamma_{0}$, denoted $\Lambda(\Gamma_{0})$, is the set of all points $x\in\partial(\Gamma,\mathcal{P})$ where there is a sequence $\{h_{n}\}$ in $\Gamma_{0}$ and $y\in\partial(\Gamma,\mathcal{P})$ such that $h_{n}(z)\rightarrow x$ for all $z\in\partial(\Gamma,\mathcal{P})-\{y\}$. Notice that the action of $\Gamma_{0}$ on $\Lambda(\Gamma_{0})$ is a convergence group action. Then $\Gamma_{0}$ is relatively quasiconvex if $\Gamma_{0}$ acts geometrically finitely on $\Lambda(\Gamma_{0})$, see [24, Defn. 6.2]. In this section we prove that infinite index, relatively quasiconvex subgroups of a relatively Anosov group have strictly lower critical exponent.
Corollary 9.1 (Corollary 1.7).
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, $\phi\in{\mathfrak{a}}_{\theta}^{*}$ and $\delta^{\phi}(\Gamma)<+\infty$. If $\Gamma_{0}$ is infinite index relatively quasiconvex subgroup of $(\Gamma,\mathcal{P})$, then
$\delta^{\phi}(\Gamma)>\delta^{\phi}(\Gamma_{0}).$
Remark 9.2.
Since we will be using results from Hruska [24] extensively in this section, we remark that the definition of relatively hyperbolic group used in [24, Defn. 3.1] is strictly weaker than our Definition 3.1. To obtain our Definition 3.1 from Hruska’s definition, one needs to further require that $\Gamma$ be non-elementary and finitely generated, and that every $P\in\mathcal{P}$ is finitely generated. By [24, Cor. 9.2], these additional conditions follow if $\Gamma$ is non-elementary and every $P\in\mathcal{P}$ is slender, i.e. any subgroup of $P$ is finitely generated.
As mentioned in the Introduction, the two results needed to deduce Corollary 9.1 from Theorems 1.1 and 1.6 are stated as Lemma 9.3 and Proposition 9.4 below. They both follow from results of Hruska [24].
Lemma 9.3.
Suppose $\Gamma$ is hyperbolic relative to $\mathcal{P}$ and $\Gamma_{0}\subset\Gamma$ is relatively quasiconvex. If $\Gamma_{0}$ has infinite index in $\Gamma$, then $\Lambda(\Gamma_{0})$ is a closed proper subset of $\partial(\Gamma,\mathcal{P})$.
Proof.
We will prove the contrapositive: if $\Lambda(\Gamma_{0})=\partial(\Gamma,\mathcal{P})$, then $\Gamma_{0}$ has finite index in $\Gamma$.
Let $X$ be a Groves–Manning cusp space for $(\Gamma,\mathcal{P})$. Given a subset $A\subset X$, we will let $\mathcal{N}_{r}(A)$ denote the closed $r$-neighborhood of $A$ in $X$. Since $X$ is Gromov hyperbolic, there exists $c>0$ such that if $\sigma_{1},\sigma_{2}:\operatorname{\mathbb{R}}\rightarrow X$ are geodesics with
$\lim_{t\rightarrow\pm\infty}\sigma_{1}(t)=\lim_{t\rightarrow\pm\infty}\sigma_{2}(t),$
then
$\sigma_{1}\subset\mathcal{N}_{c}(\sigma_{2})\quad\text{and}\quad\sigma_{2}\subset\mathcal{N}_{c}(\sigma_{1}).$
Fix a geodesic $\sigma_{0}:\operatorname{\mathbb{R}}\rightarrow X$. Then fix $r>0$ such that
$\operatorname{\mathrm{id}}\in\mathcal{N}_{r}(\sigma_{0}).$
An equivalent definition [24, Defn. 6.6] of relatively quasiconvex subgroups implies that there exists $R>0$ such that: if $s:[0,T]\rightarrow X$ is a geodesic segment with endpoints in $\Gamma_{0}$, then
$s\cap\mathcal{N}_{c+r+1}(\Gamma)\subset\mathcal{N}_{R}(\Gamma_{0}),$
see [24, Prop. 7.5 and 7.6].
Fix $g\in\Gamma$ and let $\sigma:=g\circ\sigma_{0}:\operatorname{\mathbb{R}}\rightarrow X$. Since $\Lambda(\Gamma_{0})=\partial(\Gamma,\mathcal{P})$, there exist sequences $\{h_{n}^{-}\}$, $\{h_{n}^{+}\}$ in $\Gamma_{0}$ such that
$\lim_{n\rightarrow\infty}h_{n}^{\pm}=\lim_{t\rightarrow\pm\infty}\sigma(t).$
Let $\sigma_{n}$ be a geodesic in $X$ joining $h_{n}^{-}$ to $h_{n}^{+}$. After possibly reparametrizing and passing to a subsequence we may suppose that $\sigma_{n}$ converges locally uniformly to a geodesic $\sigma_{\infty}:\operatorname{\mathbb{R}}\rightarrow X$. Then
$\sigma_{\infty}\cap\mathcal{N}_{c+r}(\Gamma)\subset\mathcal{N}_{R}(\Gamma_{0}).$
Further,
$\lim_{t\rightarrow\pm\infty}\sigma_{\infty}(t)=\lim_{n\rightarrow\infty}h_{n}^{\pm}=\lim_{t\rightarrow\pm\infty}\sigma(t),$
and so $g\in\mathcal{N}_{r}(\sigma)\subset\mathcal{N}_{c+r}(\sigma_{\infty})$. Hence,
$g\in\mathcal{N}_{R+c+r}(\Gamma_{0}).$
Since $g\in\Gamma$ was arbitrary,
$\Gamma\subset\mathcal{N}_{R+c+r}(\Gamma_{0})$
and thus $\Gamma_{0}\subset\Gamma$ has finite index. ∎
Proposition 9.4.
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$. Assume $\Gamma_{0}\subset\Gamma$ is non-elementary and relatively quasiconvex. Let $\mathcal{P}_{0}$ denote a set of representatives of the conjugacy classes in $\Gamma_{0}$ of the intersection of the peripheral subgroups of $\Gamma$ with $\Gamma_{0}$. Then:
  1. (1)
    $(\Gamma_{0},\mathcal{P}_{0})$ is relatively hyperbolic (in the sense of Definition 3.1).
  2. (2)
    $\Gamma_{0}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}_{0}$.
Proof.
Proof of (1). Hruska [24, Thm. 9.1] proved that relatively quasiconvex subgroups of relatively hyperbolic groups in his weaker sense (see Remark 9.2) are also relatively hyperbolic. Thus, it suffices to show that every $P\in\mathcal{P}_{0}$ is slender.
Lemma 9.5.
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$. If $P\in\mathcal{P}$ and $Q\subset P$ is a subgroup, then $Q$ is finitely generated.
Proof.
Using Proposition 4.5 we may assume that $Z(\mathsf{G})$ is trivial and $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$ (notice that if $p:\mathsf{G}\rightarrow\mathsf{G}^{\prime}$ is as in the proposition, then $p|_{\Gamma}$ has finite kernel and hence $Q$ is finitely generated if and only if $p(Q)$ is finitely generated). By Theorem 4.4 there exists a closed subgroup $\mathsf{H}\subset\mathsf{G}$ with finitely many components such that:
  1. (1)
    $P$ is a cocompact lattice in $\mathsf{H}$.
  2. (2)
    $\mathsf{H}=\mathsf{L}\ltimes\mathsf{U}$ where $\mathsf{L}$ is compact and $\mathsf{U}$ is the unipotent radical of $\mathsf{H}$.
Then Auslander’s theorem (see for instance [26, Thm. 11.1]) implies that $Q$ is finitely generated.
Proof of (2). By definition, $\Gamma_{0}$ is a $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup of $\mathsf{G}$. Also, by (1), $(\Gamma_{0},\mathcal{P}_{0})$ is relatively hyperbolic, and so we may identify
$\partial(\Gamma_{0},\mathcal{P}_{0})=\Lambda(\Gamma_{0})\subset\partial(\Gamma,\mathcal{P}).$
Since $\Gamma$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$, there is a $\Gamma$-equivariant homeomorphism $\xi:\partial(\Gamma,\mathcal{P})\rightarrow\Lambda_{\theta}(\Gamma)$. Thus, to show that $\Gamma_{0}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}_{0}$, it suffices to show that $\xi(\Lambda(\Gamma_{0}))=\Lambda_{\theta}(\Gamma_{0})$.
Fix any $F^{+}\in\Lambda_{\theta}(\Gamma_{0})$. Then by Proposition 2.2 there is $\{h_{n}\}$ in $\Gamma_{0}$ and $F^{-}\in\Lambda_{\theta}(\Gamma)$ such that $h_{n}(F)\rightarrow F^{+}$ for all $F\in\operatorname{\mathcal{F}}_{\theta}$ transverse to $F^{-}$. By definition, $F^{\pm}=\xi(x^{\pm})$ for some $x^{\pm}\in\partial(\Gamma,\mathcal{P})$. Also, since $\Gamma_{0}$ acts on $\Lambda(\Gamma_{0})$ as a convergence group, by passing to a subsequence, we can suppose that there exist $y^{\pm}\in\Lambda(\Gamma_{0})$ such that $h_{n}(z)\rightarrow y^{+}$ for all $z\in\Lambda(\Gamma_{0})-\{y^{-}\}$. Since $\Gamma_{0}$ is non-elementary, we may fix $z\in\Lambda(\Gamma_{0})-\{x^{-},y^{-}\}$. Then
$F^{+}=\lim_{n\rightarrow\infty}h_{n}(\xi(z))=\lim_{n\rightarrow\infty}\xi(h_{n}(z))=\xi(y^{+}).$
Since $F^{+}$ was arbitrary, it follows that $\Lambda_{\theta}(\Gamma_{0})\subset\xi(\Lambda(\Gamma_{0}))$. A very similar argument shows that $\xi(\Lambda(\Gamma_{0}))\subset\Lambda_{\theta}(\Gamma_{0})$. ∎
We may now give the proof of Corollary 9.1.
Proof of Corollary 9.1.
First, notice that if $\Gamma_{0}\subset\Gamma$ is an elementary subgroup, then it is either finite (in which case $\Lambda(\Gamma_{0})$ is empty), conjugate to a subgroup of a peripheral subgroup of $\Gamma$ (in which case $\Lambda(\Gamma_{0})$ is a single point) or virtually a cyclic group generated by a hyperbolic element (in which case $\Lambda(\Gamma_{0})$ consists of the the attracting and repelling fixed points of the hyperbolic element). If $\Gamma_{0}$ is non-elementary, then $\Lambda(\Gamma_{0})$ is perfect. See the discussion in [24, Sec. 3.1] for more details.
Case 1: Suppose $\Gamma_{0}$ is non-elementary. Let $\mathcal{P}_{0}$ denote a set of representatives of the conjugacy classes in $\Gamma_{0}$ of the intersection of the peripheral subgroups of $\Gamma$ with $\Gamma_{0}$. Then Proposition 9.4 implies that $\Gamma_{0}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}_{0}$. By Lemma 9.3, $\Lambda_{\theta}(\Gamma_{0})$ is a proper subset of $\Lambda_{\theta}(\Gamma)$. By Theorem 1.1, $Q^{\phi}_{\Gamma_{0}}$ converges at its critical exponent. So, by Theorem 1.6,
$\delta^{\phi}(\Gamma_{0})<\delta^{\phi}(\Gamma).$
Case 2: Assume #$\#\Lambda(\Gamma_{0})=2$. Then there is an infinite order hyperbolic element $\gamma\in\Gamma_{0}$ such that $\langle\gamma\rangle$ has finite index in $\Gamma_{0}$. Since $\Gamma$ acts as a convergence group on $\Lambda_{\theta}(\Gamma)$, we can label the fixed points $F^{+},F^{-}\in\Lambda_{\theta}(\Gamma)$ of $\gamma$ so that $\gamma^{n}(F)\rightarrow F^{+}$ for all $F\in\operatorname{\mathcal{F}}_{\theta}$ transverse to $F^{-}$. By Proposition 2.2, $U_{\theta}(\gamma^{-n})\rightarrow F^{-}$ and $U_{\theta}(\gamma^{n})\rightarrow F^{+}$. Then by Proposition 5.1 there exists $C>0$ such that
$\phi(\kappa_{\theta}(\gamma^{n+m}))\geq\phi(\kappa_{\theta}(\gamma^{n}))+\phi(\kappa_{\theta}(\gamma^{m}))-C$
for all $n,m\geq 1$. This estimate implies that $\delta^{\phi}(\langle\gamma\rangle)=0$, and hence that $\delta^{\phi}(\Gamma_{0})=0$.
Since $\Gamma$ is non-elementary, it contains a free subgroup of rank two and hence
$\delta^{\phi}(\Gamma_{0})=0<\delta^{\phi}(\Gamma).$
Case 3: Assume #$\#\Lambda(\Gamma_{0})=1$. Then, after conjugating, there is a peripheral subgroup $P\in\mathcal{P}$ with $\Gamma_{0}\subset P$. Then by Corollary 7.2 we have
$\delta^{\phi}(\Gamma_{0})\leq\delta^{\phi}(P)<\delta^{\phi}(\Gamma).$
Case 4: Assume #$\#\Lambda(\Gamma_{0})=0$. Then $\Gamma_{0}$ is finite and so $\delta^{\phi}(\Gamma_{0})=0$. So, as in Case 2,
$\delta^{\phi}(\Gamma_{0})=0<\delta^{\phi}(\Gamma).\qed$
10. Characterizing linear functions with finite critical exponent
In this section, we give a complete analysis of which linear functionals in $\mathfrak{a}_{\theta}^{*}$ have associated Poincaré series with finite critical exponents. This generalizes the results of Sambarino [38] for Anosov groups.
Given a subgroup $\Gamma\subset\mathsf{G}$, the $\theta$-Benoist limit cone of $\Gamma$, denoted $B_{\theta}(\Gamma)\subset\mathfrak{a}_{\theta}$, is the set of vectors $X\in\mathfrak{a}_{\theta}$ for which there exists a sequence $\{\gamma_{n}\}$ of distinct elements of $\Gamma$ and a sequence $\{r_{n}\}$ in $\mathbb{R}_{+}$ so that $r_{n}\kappa_{\theta}(\gamma_{n})$ converges to $X$.
As in Proposition 4.5, let $p:\mathsf{G}\rightarrow\mathsf{G}^{\prime}$ denote the projection map of $\mathsf{G}$ onto $\mathsf{G}^{\prime}:=\mathsf{G}/\mathsf{H}$, where $\mathsf{H}$ is the product of $Z(\mathsf{G})$ and the simple factors of $\mathsf{G}$ contained in $\operatorname{\mathsf{P}}_{\theta}$. Notice that the Benoist limit cones of a group $\Gamma\subset\mathsf{G}$ and its projection $\Gamma^{\prime}:=p(\Gamma)\subset\mathsf{G}^{\prime}$ satisfy
${\rm d}p(B_{\theta}(\Gamma))=B_{\theta^{\prime}}(\Gamma^{\prime}).$
We prove the following expanded version of Theorem 1.3.
Theorem 10.1.
Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$ and $\phi\in\mathfrak{a}_{\theta}^{*}$. The following are equivalent:
  1. (1)
    $\lim_{n\rightarrow\infty}\phi(\kappa_{\theta}(\gamma_{n}))=+\infty$ for every sequence of distinct elements $\{\gamma_{n}\}$ in $\Gamma$.
  2. (2)
    $\delta^{\phi}(\Gamma)<+\infty$.
  3. (3)
    If $x_{0}\in M^{\prime}:=\mathsf{G}^{\prime}/\operatorname{\mathsf{K}}^{\prime}$, where $\mathsf{K}^{\prime}:=p(\mathsf{K})$, then there exist constants $c\geq 1,C\geq 0$ such that
    $\frac{1}{c}{\rm d}_{M^{\prime}}(\gamma(x_{0}),x_{0})-C\leq\phi(\kappa_{\theta}(\gamma))\leq c{\rm d}_{M^{\prime}}(\gamma(x_{0}),x_{0})+C$
    for all $\gamma\in\Gamma$ (where ${\rm d}_{M^{\prime}}$ is the distance defined on $M^{\prime}$ in Section 2).
  4. (4)
    $\phi(Y)>0$ for all $Y\in B_{\theta}(\Gamma)-\{0\}$.
  5. (5)
    If $X$ is a Groves–Manning cusp space for $(\Gamma,\mathcal{P})$, then there exist constants $c,C>0$ such that
    $\phi(\kappa_{\theta}(\gamma))\geq c{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})-C$
    for all $\gamma\in\Gamma$.
Indeed, statements (1) and (2) of Theorem 1.3 are the same as statements (1) and (2) of Theorem 10.1. Furthermore, when $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$, then $\mathsf{G}^{\prime}=\mathsf{G}/Z(\mathsf{G})$. Thus, in this case $M=M^{\prime}$ and so statement (3) of Theorem 1.3 is equivalent to statement (3) of Theorem 10.1.
We will now prove Theorem 10.1. By part (3) of Proposition 4.5, we may assume that $\mathsf{G}$ has trivial center and $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$, in which case $M^{\prime}=M:=\mathsf{G}/\operatorname{\mathsf{K}}$. Then note that any one of (2), (3), (4) or (5) immediately imply (1), and (3) also immediately implies (4). Since $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$, the equivalence of (3) and (5) follows from Theorem 4.4. So it suffices to show that (3) implies (2) and (1) implies (5).
Proof of (3) $\Rightarrow$ (2): Without loss of generality we may assume that $x_{0}:=\operatorname{\mathsf{K}}\in M=\mathsf{G}/\operatorname{\mathsf{K}}$. By assumption, there exist constants $c,C>0$ such that
$\phi(\kappa_{\theta}(\gamma))\geq c{\rm d}_{M}(\gamma(x_{0}),x_{0})-C$
for all $\gamma\in\Gamma$. Then $\delta^{\phi}(\Gamma)\leq\frac{1}{c}\delta_{M}(\Gamma)$, where
$\delta_{M}(\Gamma):=\lim_{T\to\infty}\frac{\log\#\left\{\gamma\in\Gamma:{\rm d}_{M}(\gamma(x_{0}),x_{0})<T\right\}}{T}.$
Recall that the volume growth entropy of $M$ is
$h(M):=\limsup_{T\rightarrow\infty}\frac{\log{\rm Vol}_{M}(B_{T}(x_{0}))}{T}$
where ${\rm Vol}_{M}$ is the Riemannian volume on $M$ and $B_{r}(x_{0})\subset M$ is the open ball of radius $r>0$. Since $M$ has bounded sectional curvature, volume comparison theorems imply that $h(M)<+\infty$.
Fix $r_{0}>0$ and for $T>0$ let $\Gamma_{T}:=\left\{\gamma\in\Gamma:{\rm d}_{M}(\gamma(x_{0}),x_{0})<T\right\}$. Then
#$\#\Gamma_{T}=\frac{1}{{\rm Vol}_{M}(B_{r_{0}}(x_{0}))}\sum_{\gamma\in\Gamma_{T}}{\rm Vol}_{M}(B_{r_{0}}(\gamma x_{0}))\leq\frac{\#\Gamma_{2r_{0}}}{{\rm Vol}_{M}(B_{r_{0}}(x_{0}))}{\rm Vol}_{M}(B_{T+r_{0}}(x_{0})).$
Thus $\delta_{M}(\Gamma)\leq h(M)<+\infty$. ∎
The proof that (1) implies (5) is more technical, so we provide a brief outline. We first use Proposition 2.3 to provide a lower bound for $\phi\circ\kappa_{\theta}$ on peripheral subgroups, see Lemma 10.2. We then divide a geodesic joining $\operatorname{\mathrm{id}}$ to $\gamma\in\Gamma$ in the Groves–Manning cusp space $X$ into segments $\overline{\gamma_{i}\gamma_{i+1}}$ with endpoints in $\Gamma$ which either (a) have a pre-chosen size guaranteeing that $\phi(\kappa_{\theta}(\rho(\gamma_{i+1}\gamma_{i}^{-1})))$ is large enough, or (b) are at least as long as the pre-chosen size and lie entirely in a cusped portion of $X$. We then apply Proposition 5.3 to show that the image of the Cartan projections are roughly additive along the segment.
Proof of (1) $\Rightarrow$ (5): Suppose that
$\lim_{n\rightarrow\infty}\phi(\kappa_{\theta}(\gamma_{n}))=+\infty$
for every sequence of distinct elements $\{\gamma_{n}\}$ in $\Gamma$.
Fix a Groves–Manning cusp space $X$ for $(\Gamma,\mathcal{P})$. We first control the growth of peripheral elements.
Lemma 10.2.
There exist $c_{1},C_{1}>0$ such that: if $P\in\mathcal{P}$ and $\upsilon\in P$, then
$\phi(\kappa_{\theta}(\upsilon))\geq c_{1}{\rm d}_{X}(\upsilon,\operatorname{\mathrm{id}})-C_{1}.$
Proof.
Recall that ${\rm d}_{M}$ was defined so that ${\rm d}_{M}(g\operatorname{\mathsf{K}},\operatorname{\mathsf{K}})=\left\|\kappa(g)\right\|$ for all $g\in\mathsf{G}$. Then by Theorem 4.4 it suffices to find $c_{1},C_{1}>0$ such that: if $P\in\mathcal{P}$ and $\upsilon\in P$, then
$\phi(\kappa_{\theta}(\upsilon))\geq c_{1}\left\|\kappa(\upsilon)\right\|-C_{1}.$
Then, since $\mathcal{P}$ is finite, it is enough to fix $P\in\mathcal{P}$ and find constants $c_{P},C_{P}>0$ such that: if $\upsilon\in P$, then
$\phi(\kappa_{\theta}(\upsilon))\geq c_{P}\left\|\kappa(\upsilon)\right\|-C_{P}.$
By Theorem 4.4 there exists a closed subgroup $\mathsf{H}\subset\mathsf{G}$ with finitely many components such that:
  1. (1)
    $P$ is a cocompact lattice in $\mathsf{H}$.
  2. (2)
    $\mathsf{H}=\mathsf{L}\ltimes\mathsf{U}$ where $\mathsf{L}$ is compact and $\mathsf{U}$ is the unipotent radical of $\mathsf{H}$.
Let $\mathfrak{u}$ denote the Lie algebra of $\mathsf{U}$. By Proposition 2.3, $\mathsf{U}=\exp(\mathfrak{u})$. By the same proposition, for any $\alpha\in\theta$ there exist $M_{\alpha}\in\operatorname{\mathbb{N}}$, $C_{\alpha}>1$ and a positive everywhere defined rational function $R_{\alpha}:\mathfrak{u}\rightarrow\operatorname{\mathbb{R}}$ where
$\frac{1}{C_{\alpha}}R_{\alpha}(Y)^{1/M_{\alpha}}\leq e^{\omega_{\alpha}(\kappa(\exp(Y)))}\leq C_{\alpha}R_{\alpha}(Y)^{1/M_{\alpha}}$
for all $Y\in\mathfrak{u}$.
Write $\phi=\sum_{\alpha\in\theta}c_{\alpha}\omega_{\alpha}$, and define $R:=\prod_{\alpha\in\theta}R_{\alpha}^{\left|c_{\alpha}\right|/M_{\alpha}}$ and $C_{\phi}:=\prod_{\alpha\in\theta}C_{\alpha}^{\left|c_{\alpha}\right|}$. Then
$\frac{1}{C_{\phi}}R(Y)\leq e^{\phi(\kappa_{\theta}(\exp(Y)))}\leq C_{\phi}R(Y)$
for all $s\in\operatorname{\mathbb{R}}$. We proved in Lemma 7.3 that $R$ is positive and proper, so by Theorem 6.1 there exist $c_{2},\epsilon>0$ such that
$R(Y)\geq c_{2}(1+\left\|Y\right\|)^{\epsilon}$
for all $Y\in\mathfrak{u}$. By Proposition 2.3, there exist $A>0$ such that
$\left\|\kappa(\exp(Y))\right\|\leq A+A\log(1+\left\|Y\right\|)$
for all $Y\in\mathfrak{u}$. Finally, let
$M:=\max\{\left\|\kappa(\ell)\right\|:\ell\in\mathsf{L}\}.$
If $\upsilon\in P$, then $\upsilon=\ell\exp(Y)$ for some $\ell\in\mathsf{L}$ and $Y\in\mathfrak{u}$. So by Lemma 2.1,
$\phi(\kappa_{\theta}(\upsilon))$ $\geq\phi(\kappa_{\theta}(\exp(Y)))-M\left\|\phi\right\|$
$\geq\log R(Y)-M\left\|\phi\right\|-\log C_{\phi}$
$\geq\epsilon\log(1+\left\|Y\right\|)-M\left\|\phi\right\|-\log C_{\phi}+\log c_{2}$
$\geq\frac{\epsilon}{A}\left\|\kappa(\exp(Y))\right\|-\frac{\epsilon}{A}-M\left\|\phi\right\|-\log C_{\phi}+\log c_{2}$
$\geq\frac{\epsilon}{A}\left\|\kappa(\upsilon)\right\|-\frac{\epsilon}{A}M-\frac{\epsilon}{A}-M\left\|\phi\right\|-\log C_{\phi}+\log c_{2}.\qed$
By Proposition 5.3 there exists $C_{0}>0$ such that: if $f:[0,T]\rightarrow X$ is a geodesic with $f(0)=\operatorname{\mathrm{id}}$, and $f(t_{1}),f(t_{2})\in\Gamma$ for some $0\leq t_{1}\leq t_{2}\leq T$, then
$\left|\phi\big{(}\kappa_{\theta}(f(t_{2}))\big{)}-\phi\big{(}\kappa_{\theta}(f(t_{1}))\big{)}-\phi\big{(}\kappa_{\theta}(f(t_{1})^{-1}f(t_{2}))\big{)}\right|\leq C_{0}.$ (16)
By hypothesis, $\lim_{n\rightarrow\infty}\phi(\kappa_{\theta}(\gamma_{n}))=+\infty$ for every sequence of distinct elements $\{\gamma_{n}\}$ in $\Gamma$, so there exists $T_{0}>0$ such that: if $\gamma\in\Gamma$ and ${\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})\geq T_{0}$, then
$\phi(\kappa_{\theta}(\gamma))>1+C_{0}.$ (17)
Then let
$C_{2}:=\max\{\left|\phi(\kappa_{\theta}(\gamma))\right|:{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})<T_{0}\},$
let
$B:=\frac{2}{c_{1}T_{0}}\left(3C_{0}+C_{1}+2C_{2}+c_{1}T_{0}\right)>2,$
and let
$c:=\min\left\{\frac{c_{1}}{2},\frac{1}{(B+2)T_{0}}\right\}\quad\text{and}\quad C:=C_{2}+cT_{0}.$
We will show that (5) holds with $c$ and $C$ as described above.
Fix $\gamma\in\Gamma$ and let $f:[0,T]\rightarrow X$ be a geodesic with $f(0)=\operatorname{\mathrm{id}}$ and $f(T)=\gamma$. If $T<T_{0}$, then
$\phi(\kappa_{\theta}(\gamma))\geq-C_{2}\geq c{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})-C.$
If $T\geq T_{0}$, fix a partition
$0=t_{0}<t_{1}<\dots<t_{m}=T$
with the following properties:
  1. (1)
    $t_{n+1}-t_{n}\geq T_{0}$ for $n=0,\dots,m-1$,
  2. (2)
    $\gamma_{n}:=f(t_{n})\in\Gamma$ for $n=0,\dots,m$ and
  3. (3)
    if $0=s_{0}<s_{1}<\dots<s_{m^{\prime}}=T$ is another partition with the first two properties, then $m^{\prime}\leq m$.
Then by Equation (16),
$\phi(\kappa_{\theta}(\gamma))=\sum_{n=0}^{m-1}\Big{(}\phi(\kappa_{\theta}(\gamma_{n+1}))-\phi(\kappa_{\theta}(\gamma_{n}))\Big{)}\geq\sum_{n=0}^{m-1}\Big{(}\phi(\kappa_{\theta}(\gamma_{n}^{-1}\gamma_{n+1}))-C_{0}\Big{)}.$
Thus to complete the proof it suffices to verify that
$\phi(\kappa_{\theta}(\gamma_{n}^{-1}\gamma_{n+1}))\geq c(t_{n+1}-t_{n})+C_{0}$ (18)
for each $n=0,\dots,m-1$. Indeed, if this were the case, then
$\phi(\kappa_{\theta}(\gamma))\geq c\sum_{n=0}^{m-1}(t_{n+1}-t_{n})=cT\geq c{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})-C.$
We will now prove Equation (18). Fix $n\in\{0,\dots,m-1\}$. If $t_{n+1}-t_{n}<(B+2)T_{0}$, then Equation (17) implies that
$\phi(\kappa_{\theta}(\gamma_{n}^{-1}\gamma_{n+1}))\geq 1+C_{0}\geq c(t_{n+1}-t_{n})+C_{0}.$
If $t_{n+1}-t_{n}>(B+2)T_{0}$, then by the maximality of the partition and the fact that $B+2>4$, there exist $P\in\mathcal{P}$, $\eta\in\Gamma$ and $a,b\in[t_{n},t_{n+1}]$ such that:
  1. (1)
    $a\in[t_{n},t_{n}+T_{0})$, $b\in(t_{n+1}-T_{0},t_{n+1}]$ (hence $a<b$),
  2. (2)
    $f(a),f(b)\in\eta P$, and
  3. (3)
    $f|_{[a,b]}$ is contained in the combinatorial horoball associated to $\eta P$.
Then by applying Equation (16) to the geodesic $t\mapsto\gamma_{n}^{-1}f(t)$,
$\phi(\kappa_{\theta}(\gamma_{n}^{-1}\gamma_{n+1}))$ $\geq\phi(\kappa_{\theta}(\gamma_{n}^{-1}f(a)))+\phi(\kappa_{\theta}(f(a)^{-1}f(b)))+\phi(\kappa_{\theta}(f(b)^{-1}\gamma_{n+1}))-2C_{0}$
$\geq\phi(\kappa_{\theta}(f(a)^{-1}f(b)))-2C_{0}-2C_{2},$
where the last equality holds by the definition of $C_{2}$. Since $f(a)^{-1}f(b)\in P$ and
${\rm d}_{X}(f(a)^{-1}f(b),\operatorname{\mathrm{id}})={\rm d}_{X}(f(a),f(b))>BT_{0},$
by Lemma 10.2, we then have
$\phi(\kappa_{\theta}(\gamma_{n}^{-1}\gamma_{n+1}))$ $\geq c_{1}{\rm d}_{X}(f(a)^{-1}f(b),\operatorname{\mathrm{id}})-2C_{0}-C_{1}-2C_{2}$
$\geq\frac{c_{1}}{2}{\rm d}_{X}(f(a),f(b))+C_{0}+c_{1}T_{0}$
$\geq c(t_{n}-t_{n-1})+C_{0}.$
This completes the proof. ∎
Appendix A Proof of Proposition 2.3
In this appendix, we prove Proposition 2.3. We start with an observation about the linear case.
Lemma A.1.
If $\mathsf{U}\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ is a connected unipotent group with Lie algebra $\mathfrak{u}$, then there exist $C_{0}>1$ and a positive polynomial $P:\mathfrak{u}\rightarrow\operatorname{\mathbb{R}}$ such that
$\frac{1}{C_{0}}P(Y)^{1/2}\leq\sigma_{1}(e^{Y})\leq C_{0}P(Y)^{1/2}$
for all $Y\in\mathfrak{u}$.
Proof.
Define $P:\mathfrak{u}\rightarrow\operatorname{\mathbb{R}}$ by
$P(Y)=\sum_{1\leq i,j\leq d}\left[e^{Y}\right]_{i,j}^{2}.$
Since $\mathfrak{u}$ is nilpotent, see  [7, Section 4.8], $P$ is a polynomial. Observe that the Euclidean norm $\left\|\cdot\right\|_{\rm Euc}:\mathrm{End}(\operatorname{\mathbb{R}}^{d})\to\operatorname{\mathbb{R}}$ and the first singular value $\sigma_{1}:\mathrm{End}(\operatorname{\mathbb{R}}^{d})\to\operatorname{\mathbb{R}}$ are both norms on the vector space $\mathrm{End}(\operatorname{\mathbb{R}}^{d})$, so there exists $C_{0}>1$ such that
$\frac{1}{C_{0}}\leq\frac{\sigma_{1}(X)}{\left\|X\right\|_{\rm Euc}}\leq C_{0}$
for all $X\in\mathrm{End}(\operatorname{\mathbb{R}}^{d})$. Since $P(Y)^{1/2}=\left\|e^{Y}\right\|_{\rm Euc}$ for all $Y\in\mathfrak{u}$, the lemma follows. ∎
Proof of Proposition 2.3.
Proof of (1). Since $Z(\mathsf{G})$ is trivial, $\operatorname{\mathsf{U}}$ is isomorphic to ${\rm Ad}(\operatorname{\mathsf{U}})$. Since each element of ${\rm Ad}(\operatorname{\mathsf{U}})$ is unipotent, the matrix logarithm
$\log(A)=\sum_{n=1}^{\infty}(-1)^{n+1}\frac{(A-\operatorname{\mathrm{id}})^{n}}{n}$
is well defined on ${\rm Ad}(\operatorname{\mathsf{U}})$. So the exponential map of $\mathsf{SL}(\mathfrak{g})$ induces a diffeomorphism ${\rm ad}(\mathfrak{u})\rightarrow{\rm Ad}(\mathsf{U})$, which implies that the exponential map of $\mathsf{G}$ induces a diffeomorphism $\mathfrak{u}\rightarrow\mathsf{U}$.
Proof of (2). See for instance [39, Prop. 10.14].
Proof of (3). See for instance [44, Prop. 3.4.2].
Proof of (4). Let $d:=\dim\mathfrak{g}$ and fix a linear identification $\mathfrak{g}=\operatorname{\mathbb{R}}^{d}$. This induces identifications $\mathsf{GL}(\mathfrak{g})=\mathsf{GL}(d,\operatorname{\mathbb{R}})$ and $\mathfrak{sl}(\mathfrak{g})=\mathfrak{sl}(d,\operatorname{\mathbb{R}})$. Using the root space decomposition, we can pick our identification so that ${\rm ad}(\mathfrak{a})$ is a subgroup of the diagonal matrices in $\mathsf{GL}(\mathfrak{g})=\mathsf{GL}(d,\operatorname{\mathbb{R}})$
Since $Z(\mathsf{G})$ is trivial, ${\rm ad}:\mathfrak{g}\rightarrow\mathfrak{sl}(\mathfrak{g})$ is injective, so the map
$\sigma_{1}\circ{\rm ad}:{\mathfrak{g}}\to\operatorname{\mathbb{R}}$
is a norm on $\mathfrak{g}$. Hence there exists $C_{1}>1$ such that
$\left\|X\right\|\leq C_{1}\sigma_{1}({\rm ad}(X))$
for all $X\in\mathfrak{g}$. Then, since ${\rm ad}(\mathfrak{a})$ is a subgroup of the diagonal matrices,
$\left\|Y\right\|\leq C_{1}\sigma_{1}({\rm ad}(Y))=C_{1}\log\sigma_{1}({\rm Ad}(e^{Y}))$
for all $Y\in\mathfrak{a}$. Hence, by the $\mathsf{KAK}$-decomposition,
$\left\|\kappa(g)\right\|\leq C_{1}\log\sigma_{1}({\rm Ad}(e^{\kappa(g)}))\leq C_{1}C_{2}^{2}\log\sigma_{1}({\rm Ad}(g))$
for all $g\in\mathsf{G}$, where
$C_{2}:=\max_{k\in\operatorname{\mathsf{K}}}\,\sigma_{1}({\rm Ad}(k)).$
So by Lemma A.1 and part (3),
$\left\|\kappa(e^{Y})\right\|$ $\leq C_{1}C_{2}^{2}\log\sigma_{1}(e^{{\rm ad}(Y)})\leq C_{1}C_{2}^{2}\log C_{0}+\frac{C_{1}C_{2}^{2}}{2}\log P({\rm ad}(Y))$
for all $Y\in\mathfrak{u}$. Thus, to prove (4), it suffices to show that there is some $A,a>0$ such that
$P({\rm ad}(Y))\leq A(1+\left\|Y\right\|)^{a}$ (19)
for all $Y\in{\mathfrak{u}}$.
Again, since $Z(\mathsf{G})$ is trivial, the map $\left\|\cdot\right\|^{\prime}:\mathfrak{g}\to\operatorname{\mathbb{R}}$ given by
$\left\|Y\right\|^{\prime}:=\max_{1\leq i,j\leq d}\left|[{\rm ad}(Y)]_{i,j}\right|.$
[…]
$P({\rm ad}(Y))\leq Q(\left\|Y\right\|^{\prime})$
for all $Y\in\mathfrak{u}$. Observe that there is some $B,b>0$ such that
$Q(\left\|Y\right\|^{\prime})\leq B(1+\left\|Y\right\|^{\prime})^{b}$
for all $Y\in\mathfrak{u}$, so Equation (19) holds. So (4) follows.
Proof of (5). For $\psi\in\sum_{\alpha\in\Delta}\operatorname{\mathbb{Z}}_{\geq 0}\cdot\omega_{\alpha}$ let $\chi_{\psi}:=\psi+\sum_{\alpha\in\Delta}\omega_{\alpha}$. By Proposition 4.6, for each such $\psi$ there exist $d_{\psi},N_{\psi}\in\operatorname{\mathbb{N}}$ and a irreducible representation $\Phi_{\psi}:\mathsf{G}\rightarrow\operatorname{\mathsf{SL}}(d_{\psi},\operatorname{\mathbb{R}})$ such that
$e^{\chi_{\psi}(\kappa(g))}=\sigma_{1}(\Phi_{\psi}(g))^{1/N_{\psi}}$
for all $g\in\mathsf{G}$. By part (3), $\Phi_{\psi}(\mathsf{U})\subset\operatorname{\mathsf{SL}}(d_{\psi},\operatorname{\mathbb{R}})$ is unipotent with Lie algebra ${\rm d}\Phi_{\psi}(\mathfrak{u})$. So by Lemma A.1 there exist $A_{\psi}>1$ and a positive polynomial $P_{\psi}:\mathfrak{u}\rightarrow\operatorname{\mathbb{R}}$ such that
$\frac{1}{A_{\psi}}P_{\psi}(Y)^{1/(2N_{\psi})}\leq e^{\chi_{\psi}(\kappa(e^{Y}))}\leq A_{\psi}P_{\psi}(Y)^{1/(2N_{\psi})}$ (20)
for all $Y\in\mathfrak{u}$. Since $\omega_{\alpha}=\chi_{\omega_{\alpha}}-\chi_{0}$, (5) follows from Equation (20) with
$R_{\alpha}(Y):=\frac{P_{\omega_{\alpha}}(Y)^{N_{0}}}{P_{0}(Y)^{N_{\omega_{\alpha}}}},$
$M_{\alpha}:=2N_{0}N_{\omega_{\alpha}}$ and $C_{\alpha}:=A_{\omega_{\alpha}}A_{0}$. ∎
Appendix B Proof of Theorem 4.4
In this appendix we prove Theorem 4.4. As mentioned before, Theorem 4.4 was established in [43] in the special case when $\mathsf{G}=\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$. In the following argument we use Proposition 4.6 to reduce to this special case.
Theorem B.1.
Assume $Z(\mathsf{G})$ is trivial and $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$. Suppose $\Gamma\subset\mathsf{G}$ is a $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup relative to $\mathcal{P}$.
  1. (1)
    If $X$ is a Groves–Manning cusp space for $(\Gamma,\mathcal{P})$ and $M:=\mathsf{G}/\operatorname{\mathsf{K}}$ is a Riemannian symmetric space associated to $\mathsf{G}$, then there exist $c>1$, $C>0$ such that
    $\frac{1}{c}{\rm d}_{M}(\gamma\operatorname{\mathsf{K}},\operatorname{\mathsf{K}})-C\leq{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})\leq c{\rm d}_{M}(\gamma\operatorname{\mathsf{K}},\operatorname{\mathsf{K}})+C$
    for all $\gamma\in\Gamma$.
  2. (2)
    If $P\in\mathcal{P}$, then $P$ is a cocompact lattice in a closed Lie group $\mathsf{H}$ with finitely many components. Moreover,
    1. (a)
      $\mathsf{H}=\mathsf{L}\ltimes\mathsf{U}$ where $\mathsf{L}$ is compact and $\mathsf{U}$ is the unipotent radical of $\mathsf{H}$.
    2. (b)
      $\mathsf{H}^{0}=\mathsf{L}^{0}\times\mathsf{U}$ and $\mathsf{L}^{0}$ is Abelian.
Fix $\mathsf{G}$, $\operatorname{\mathsf{P}}_{\theta}$, $\Gamma$, and $\mathcal{P}$ satisfy the assumptions of Theorem 4.4. Let $\chi:=\sum_{\alpha\in\theta}\omega_{\alpha}$, and let $N\in\operatorname{\mathbb{N}}$, $\Phi:\mathsf{G}\rightarrow\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ and $\xi:\operatorname{\mathcal{F}}_{\theta}\rightarrow\operatorname{\mathcal{F}}_{1,d-1}(\operatorname{\mathbb{R}}^{d})$ satisfy Proposition 4.6 for $\chi$. Let $\mathsf{G}_{\star}$ denote the Zariski closure of $\Phi(\mathsf{G})$ in $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$.
Lemma B.2.
$\Phi$ is injective, $\mathsf{G}_{\star}$ is semisimple and $\Phi(\mathsf{G})=\mathsf{G}_{\star}^{0}$.
Proof.
Since $\mathsf{G}$ is semisimple, $\ker\Phi$ is either discrete or contains a simple factor of $\mathsf{G}$. Since $\xi:\operatorname{\mathcal{F}}_{\theta}\rightarrow\operatorname{\mathcal{F}}_{1,d-1}(\operatorname{\mathbb{R}}^{d})$ is a $\Phi$-equivariant embedding, $\ker\Phi$ must act trivially on $\operatorname{\mathcal{F}}_{\theta}$. So $\ker\Phi\subset\operatorname{\mathsf{P}}_{\theta}$. By assumption $\operatorname{\mathsf{P}}_{\theta}$ contains no simple factors of $\mathsf{G}$, so $\ker\Phi$ is discrete. Since $\mathsf{G}$ is connected, for every $g\in\mathsf{G}$, there is a continuous path in $\mathsf{G}$ connecting $\operatorname{\mathrm{id}}$ and $g$. Since $\ker\Phi$ is normal, this implies that for any $h\in\ker\Phi$, there is a continuous path in $\ker\Phi$ between $h$ and $ghg^{-1}$. The discreteness of $\ker\Phi$ then implies that $ghg^{-1}=h$. Since both $g$ and $h$ are arbitrary, we see that $\ker\Phi$ is contained in the center of $\mathsf{G}$, and so $\ker\Phi$ is trivial.
By construction $\Phi(\mathsf{G})\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ is irreducible and contains a proximal element and hence $\mathsf{G}_{\star}$ is a semisimple Lie group by  [11, Lem. 2.19].
Since ${\rm d}\Phi(\mathfrak{g})$ is the Lie algebra of $\Phi(\mathsf{G})$,
${\rm Ad}(h){\rm d}\Phi(\mathfrak{g})={\rm d}\Phi(\mathfrak{g})$
for all $h\in\Phi(\mathsf{G})$. So
${\rm Ad}(h){\rm d}\Phi(\mathfrak{g})={\rm d}\Phi(\mathfrak{g})$
for all $h\in\mathsf{G}_{\star}$. Since $\mathsf{G}$ is connected, $\Phi(\mathsf{G})$ is a connected normal subgroup of $\mathsf{G}_{\star}^{0}$, and thus is an almost direct product of simple factors of $\mathsf{G}_{\star}^{0}$.
Suppose for contradiction that there is a simple factor $\mathsf{H}\subset\mathsf{G}^{0}_{\star}$ that does not lie in $\Phi(\mathsf{G})$. Since $\Phi$ is irreducible and $\mathsf{H}$ commutes with $\Phi(\mathsf{G})$, we may apply Schur’s lemma to deduce that $\mathsf{H}\cong\operatorname{\mathbb{R}}$, which is impossible since $\mathsf{G}_{\star}^{0}$ is semisimple. Thus, $\Phi(\mathsf{G})=\mathsf{G}_{\star}^{0}$. ∎
Recall that ${\rm d}_{M}$ satisfies ${\rm d}_{M}(g\operatorname{\mathsf{K}},\operatorname{\mathsf{K}})=\left\|\kappa(g)\right\|$ for all $g\in\mathsf{G}$. So to prove part (1) it suffices to prove the following.
Lemma B.3.
If $X$ is a Groves–Manning cusp space for $(\Gamma,\mathcal{P})$, then there exist $c>1$, $C>0$ such that
$\frac{1}{c}\left\|\kappa(\gamma)\right\|-C\leq{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})\leq c\left\|\kappa(\gamma)\right\|+C$
for all $\gamma\in\Gamma$.
Proof.
Let $N:=\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})/\mathsf{SO}(d)$ be the symmetric space associated to $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ and let $x_{0}:=\mathsf{SO}(d)\in N$. Since $\Phi(\Gamma)$ is $\operatorname{\mathsf{P}}_{1,d-1}$-Anosov relative to $\mathcal{P}$ and $\ker\Phi$ is trivial, by [43, Thm. 1.7], there exist $c_{1}>1$, $C_{1}>0$ such that
$\frac{1}{c_{1}}{\rm d}_{N}(\Phi(\gamma)x_{0},x_{0})-C_{1}\leq{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})\leq c_{1}{\rm d}_{N}(\Phi(\gamma)x_{0},x_{0})+C_{1}$
for all $\gamma\in\Gamma$. Then there exist $c_{2}>1$ such that
$\frac{1}{c_{2}}\log\sigma_{1}(\Phi(\gamma))-C_{1}\leq{\rm d}_{X}(\gamma,\operatorname{\mathrm{id}})\leq c_{2}\log\sigma_{1}(\Phi(\gamma))+C_{1}$ (21)
for all $\gamma\in\Gamma$.
Lemma B.2 implies that ${\rm d}\Phi({\mathfrak{a}})$ is a Cartan subspace of the Lie algebra of $\mathsf{G}_{\star}$. By [34, Thm. 7], ${\rm d}\Phi({\mathfrak{a}})$ is conjugate to a subspace of the symmetric matrices in $\mathfrak{sl}(d,\operatorname{\mathbb{R}})$, which in turn implies that ${\rm d}\Phi({\mathfrak{a}})$ is conjugate to a subspace of the diagonal matrices. So there exists $c_{3}>1$ such that
$\frac{1}{c_{3}}e^{\sigma_{1}({\rm d}\Phi(X))}\leq\sigma_{1}(e^{{\rm d}\Phi(X)})\leq c_{3}e^{\sigma_{1}({\rm d}\Phi(X))}$
for all $X\in\mathfrak{a}$. Since $\Phi$ is injective, so is ${\rm d}\Phi$. Hence there exists $c_{4}>1$ such that
$\frac{1}{c_{4}}\sigma_{1}({\rm d}\Phi(X))\leq\left\|X\right\|\leq c_{4}\sigma_{1}({\rm d}\Phi(X))$
for all $X\in\mathfrak{g}$. Finally, since $\operatorname{\mathsf{K}}$ is compact,
$D:=\max_{k\in\operatorname{\mathsf{K}}}\sigma_{1}(\Phi(k))$
is finite.
Now if $g\in\mathsf{G}$, then by the $\mathsf{KAK}$-decomposition,
$\log\sigma_{1}(\Phi(g))$ $\leq 2\log D+\log\sigma_{1}\left(e^{{\rm d}\Phi(\kappa(g))}\right)$
$\leq 2\log D+\log c_{3}+\sigma_{1}({\rm d}\Phi(\kappa(g)))$
$\leq 2\log D+\log c_{3}+c_{4}\left\|\kappa(g)\right\|$
and likewise
$\log\sigma_{1}(\Phi(g))\geq-2\log D-\log c_{3}+\frac{1}{c_{4}}\left\|\kappa(g)\right\|.$
Combining these estimates with Equation (21) completes the proof. ∎
Lemma B.4.
If $P\in\mathcal{P}$, then $P$ is a cocompact lattice in a closed Lie subgroup $\mathsf{H}\subset\mathsf{G}$ with finitely many components. Moreover,
  1. (1)
    $\mathsf{H}=\mathsf{L}\ltimes\mathsf{U}$ where $\mathsf{L}$ is compact and $\mathsf{U}$ is the unipotent radical of $\mathsf{H}$.
  2. (2)
    $\mathsf{H}^{0}=\mathsf{L}^{0}\times\mathsf{U}$ and $\mathsf{L}^{0}$ is Abelian.
Proof.
Let $\mathsf{H}_{\star}$ denote the Zariski closure of $\Phi(P)$ in $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$. Since $\Phi(\Gamma)$ is $\operatorname{\mathsf{P}}_{1,d-1}$-Anosov relative to $\mathcal{P}$,  [43, Prop. 4.2 and Thm. 8.1] imply that $\Phi(P)$ is a cocompact lattice in $\mathsf{H}_{\star}$. Moreover
  1. (1)
    $\mathsf{H}_{\star}=\mathsf{L}_{\star}\ltimes\mathsf{U}_{\star}$ where $\mathsf{L}_{\star}$ is compact and $\mathsf{U}_{\star}$ is the unipotent radical of $\mathsf{H}_{\star}$.
  2. (2)
    $\mathsf{H}^{0}_{\star}=\mathsf{L}^{0}_{\star}\times\mathsf{U}_{\star}$ and $\mathsf{L}^{0}_{\star}$ is Abelian.
Then $\mathsf{U}:=\Phi^{-1}(\mathsf{U}_{\star})$, $\mathsf{L}:=\Phi^{-1}(\mathsf{L}_{\star}\cap\mathsf{G}_{\star}^{0})$ and $\mathsf{H}:=\Phi^{-1}(\mathsf{H}_{\star}\cap\mathsf{G}_{\star}^{0})$ satisfy the lemma. ∎